### v5

# Topological rigidity of small RCD(K,N) spaces with maximal rank

Abstract.
For a polycyclic group $\Lambda$, ${\text{rank}}(\Lambda)$ is defined as the number of $\mathbb{Z}$ factors in a polycyclic decomposition of $\Lambda$. For a finitely generated group $G$, ${\text{rank}}(G)$ is defined as the infimum of ${\text{rank}}(\Lambda)$ among finite index polycyclic subgroups $\Lambda\leq G$.
For a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ with ${\text{diam}}(X)\leq\varepsilon(K,N)$, the rank of $\pi_{1}(X)$ is at most $N$. We show that in case of equality, $X$ is homeomorphic to an infranilmanifold, generalizing a result by Kapovitch–Wilking to the non-smooth setting.
1. Introduction
A classical result of Gromov [30], refined by Ruh [63], asserts that a closed $N$-dimensional Riemannian manifold $X$ with
$\hskip 22.76228pt|\sec(X)|\leq K,\hskip 85.35826pt{\text{diam}}(X)\leq\varepsilon(K,N),$
is diffeomorphic to an infranilmanifold. Recall that an infranilmanifold is a compact smooth manifold of the form $G/\Gamma$ with $G$ a simply-connected nilpotent Lie group and $\Gamma$ a torsion-free discrete subgroup of the semi-direct product of $G$ with a compact group of automorphisms of $G$. If $\Gamma\leq G$, then the infranilmanifold is called a nilmanifold.
This full topological control is lost if one works with only a Ricci curvature bound; this is illustrated by the fact that the topology of Ricci flat manifolds is still quite mysterious. However, the fundamental group is very well understood. By the work of Kapovitch–Wilking [45], if a closed $N$-dimensional Riemannian manifold $X$ satisfies
$\hskip 25.6073pt\text{Ric}(X)\geq K,\hskip 85.35826pt{\text{diam}}(X)\leq\varepsilon(K,N),$
then $\pi_{1}(X)$ contains a subgroup $\Lambda$ of index $\pi_{1}(X):\Lambda]\leq C(K,N)$ generated by elements $\{u_{1},\ldots,u_{N}\}\subset\Lambda$ with $u_{i},u_{j}]\in\langle u_{j+1},\ldots,u_{N}\rangle$ for each $i,j\in\{1,\ldots,N\}$. Moreover, if
$\langle u_{j},\ldots,u_{N}\rangle/\langle u_{j+1},\ldots,u_{N}\rangle\cong\mathbb{Z}\text{ for all }j\in\{1,\ldots,N\},$
then $X$ is homeomorphic to an infranilmanifold. Our main result is a generalization of this maximal rank topological rigidity to the non-smooth setting. We refer the reader to Section 3.9 for the definition of rank of a finitely generated group.
Theorem A.
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality, $X$ is homeomorphic to an infranilmanifold of dimension $N$.
Theorem A can be regarded as a non-abelian analogue of the following result [52, Theorem 1.2.3]. Recall that for a path connected topological space $X$, its first Betti number $\text{b}_{1}(X)$ can be defined as the supremum of $k$ for which there is a surjective morphism $\pi_{1}(X)\to\mathbb{Z}^{k}$.
[…]
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then $\text{b}_{1}(X)\leq N$, and in case of equality, $X$ is bi-Hölder homeomorphic to a flat torus of dimension $N$.
Previoulsy, Cheeger–Colding established that under the hypotheses of Theorem 1, if $X$ is a smooth Riemannian manifold, then it is diffeomorphic to a flat torus [12, Theorem A.1.13]. As a second goal of this paper, we provide some missing details in the proofs of [52, Theorem 1.2.3] and [12, Theorem A.1.13]. We first indicate where the missing details are. By contradiction, one assumes there is a sequence of $\text{RCD}(K,N)$ spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ with
${\text{diam}}(X_{i})\to 0,\hskip 85.35826pt\text{b}_{1}(X_{i})=N,$
but $X_{i}$ not bi-Hölder homeomorphic to a flat torus. In [52], a sequence of finite sheeted covers $X_{i}^{\prime}\to X_{i}$ is constructed in such a way that $X_{i}^{\prime}$ converges in the Gromov–Hausdorff sense to the flat torus $\mathbb{T}^{N}$. From here, the Cheeger–Colding version of the Reifenberg Theorem [12, Appendix A] implies that $X_{i}^{\prime}$ is bi-Hölder homeomorphic to $\mathbb{T}^{N}$ for $i$ large enough, and it is deduced using topological arguments that the spaces $X_{i}$ are homeomorphic to $\mathbb{T}^{N}$.
Their argument concludes here, but the proof is not finished, since a bi-Hölder homeomorphism $X_{i}^{\prime}\to\mathbb{T}^{N}$ doesn’t necessarily descend to a bi-Hölder homeomorphism between $X_{i}$ and a flat torus. Moreover, if $N\geq 5$, there exists exotic tori, i.e., smooth manifolds homeomorphic to the standard $N$-torus but not diffeomorphic to it [69, Section 15.A], and any such exotic tori admits a finite-sheeted covering diffeomorphic to the standard $N$-torus [22, Theorem A.1].
We address this with the following result, which is a modification of the canonical homeomorphism constructed by Honda–Peng in [35, Theorem 6.3] (see the end of Section 6 for further details). We refer the reader to Section 3.4 for the definition of measured Gromov–Hausdorff convergence.
Theorem B.
For each $K\in\mathbb{R}$, $N\geq 1$, $\varepsilon>0$, there is $\delta>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is a compact $\text{RCD}(K,N)$ space $\delta$-close to $\mathbb{T}^{N}$ in the measured Gromov–Hausdorff sense, and there is an abelian group $G\leq\text{Iso}(X)$ with ${\text{diam}}(X/G)<\delta$, then there is a free isometric action of $G$ on $\mathbb{T}^{N}$ and a $G$-equivariant homeomorphism $F:X\to\mathbb{T}^{N}$ with
(1.1) $(1-\varepsilon)\,\mathsf{d}(x,y)^{1+\varepsilon}\leq\mathsf{d}(F(x),F(y))\leq(1+\varepsilon)\,\mathsf{d}(x,y)\,\text{ for all }x,y\in X.$
Furthermore, if $X$ is a smooth Riemannian manifold, the map $F$ can be taken to be a diffeomorphism.
In the setting of smooth Riemannian manifolds, Huang obtained a similar result for more general limit spaces and groups [36, Corollary 1.9], which can be used in an analogous way to fill the gap in the proof of [12, Theorem A.1.13].
1.1. Strategy and Outline
The Borel conjecture asserts that any two closed aspherical manifolds with isomorphic fundamental group are homeomorphic [62]. Since this conjecture is known to be true for manifolds with virtually nilpotent fundamental groups (see Theorem 4), and the Margulis Lemma holds for $\text{RCD}(K,N)$ spaces (see [17]), then the proof of Theorem A consists mainly of showing that if ${\text{diam}}(X)\leq\varepsilon$ and ${\text{rank}}(\pi_{1}(X))=N$, then $X$ is a closed aspherical topological manifold. This approach is also used in [45] for the proof of the smooth case. However, our proof is significantly different and builds upon the work of Breuillard–Green–Tao on the structure of approximate groups [7].
Theorem B is proven by contradiction. Assuming there is a sequence of $\text{RCD}(K,N)$ spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ and groups $G_{i}\leq\text{Iso}(X_{i})$ that contradict the statement of the theorem, we show that for $i$ large enough, using eigenfunctions of the Laplacian on $X_{i}$, one can find representations $G_{i}\to O(2N)$ and equivariant maps $\Phi_{i}:X_{i}\to\mathbb{R}^{2N}$ such that $\mathbb{T}^{N}\subset\mathbb{R}^{2N}$ is $G_{i}$-invariant and the maps $\Phi_{i}$ satisfy the conditions of [35, Theorem 1.2.2], which we apply to deduce (1.1).
In Section 2 we discuss related results and conjectures. In Section 3 we cover the background material we need for our main results. In Section 4 we prove the technical topological tools required for our main theorems. In Section 5 we prove Theorem A. In Section 6 we prove Theorem B and fill in the details in the proof of Theorem 1.
2. Related problems and results
Under the conditions of Theorem A, if $X$ is a smooth $N$-dimensional Riemannian manifold with ${\text{rank}}(\pi_{1}(X))=N$, then it follows from [55, Proposition 5.4] and the work of Huang–Kong–Rong–Xu [37] that $X$ is diffeomorphic to an infranilmanifold (this was also proven by Naber–Zhang [55] assuming an upper Ricci curvature bound). We note, however, that their proof relies on the Ricci flow, so their techniques do not extend to the non-smooth setting. Later Rong [61] gave a more geometric proof inspired by Cheeger–Fukaya–Gromov collapsing theory [14], but some tools used there still rely on the smooth structure. In view of this discussion and Theorem 1, it would be interesting to know if the conclusion of Theorem A can be improved to a bi-Hölder homeomorphism.
[…]
For each $K\in\mathbb{R},$ $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space with ${\text{diam}}(X)\leq\varepsilon$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality, $X$ is bi-Hölder homeomorphic to an infranilmanifold of dimension $N$.
The following proposition shows that the approach of constructing finite-sheeted covers $X_{i}^{\prime}\to X_{i}$ that don’t collapse fails when dealing with Conjecture 2 because in general, fundamental groups of infranilmanifolds are not virtually abelian.
[…]
Let $X_{i}$ be a sequence of closed $N$-dimensional Riemannian manifolds with $\text{Ric}(X_{i})\geq-\frac{1}{i}$ and $\pi_{1}(X_{i})$ torsion-free for all $i$. If there is a sequence of finite-sheeted covers $X_{i}^{\prime}\to X_{i}$ with
${\text{diam}}(X_{i}^{\prime})\leq D,\hskip 85.35826pt\mathcal{H}^{N}(X_{i}^{\prime})\geq\nu$
for some $D,\nu>0$, then $\pi_{1}(X_{i})$ is virtually abelian for large $i$.
Proof.
Arguing by contradiction, after passing to a subsequence, we can assume $\pi_{1}(X_{i})$ is not virtually abelian for any $i$. Let $\tilde{X}_{i}$ be the sequence of universal covers, $p_{i}\in X_{i}^{\prime}$, and $\tilde{p}_{i}\in\tilde{X}_{i}$ in the preimage of $p_{i}$. By Theorems 3.3 and 3.3, after further passing to a subsequence, the triples $(\tilde{X}_{i},\tilde{p}_{i},\pi_{1}(X_{i}^{\prime}))$ converge in the equivariant Gromov–Hausdorff sense to a triple $(Y,\tilde{p},G)$, while $X_{i}^{\prime}$ converges to $Y/G$.
Since the sequence $X_{i}^{\prime}$ doesn’t collapse, neither does the sequence $\tilde{X}_{i}$ and the limit group $G$ is discrete so by the work of Pan–Rong [57, Theorem 2.17] the groups $\pi_{1}(X_{i}^{\prime})$ are isomorphic to $G$ for large $i$. By the work of Pan–Wang [58, Theorem 1.5], $\pi_{1}(Y/G)$ is the quotient of $G$ by a subgroup generated by torsion elements, but since $G$ is torsion-free such subgroup must be trivial and $\pi_{1}(Y/G)=G$.
Finally, $Y/G$ equipped with $\mathcal{H}^{N}$ is a compact $\text{RCD}(0,N)$ space so its fundamental group is virtually abelian, hence so is $\pi_{1}(X_{i})$ for large $i$; a contradiction. ∎
If $\sec(X_{i})\geq-1/i$, one could remove from Proposition 2 the hypothesis that the groups $\pi_{1}(X_{i})$ are torsion-free since one could use Perelman stability theorem [42, 59] to deduce that $X_{i}^{\prime}$ and $Y/G$ are homeomorphic for large $i$. Example 2 below illustrates the intuition behind Proposition 2: in order to obtain finite-sheeted covers with controlled volume, one must “unwind” the manifold in all directions, which can lead to the diameter growing to infiniy if the groups are not abelian.
[…]
Consider the $3$-dimensional Heisenberg groups
$H^{3}(\mathbb{R})=\left\{\left.\begin{bmatrix}1&x&z\\ 0&1&y\\ 0&0&1\end{bmatrix}\right|x,y,z\in\mathbb{R}\right\},\quad H^{3}(\mathbb{Z})=\left\{\left.\begin{bmatrix}1&a&c\\ 0&1&b\\ 0&0&1\end{bmatrix}\right|a,b,c\in\mathbb{Z}\right\}.$
Let $\{\text{v}_{1},\,\text{v}_{2},\,\text{v}_{3}\}$ be the basis of the Lie algebra of $H^{3}(\mathbb{R})$ given by
$\text{v}_{1}=\begin{bmatrix}0&1&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\hskip 56.9055pt\text{v}_{2}=\begin{bmatrix}0&0&0\\ 0&0&1\\ 0&0&0\end{bmatrix},\hskip 56.9055pt\text{v}_{3}=\begin{bmatrix}0&0&1\\ 0&0&0\\ 0&0&0\end{bmatrix}$.
For $i\in\mathbb{N}$, consider the left invariant Riemannian metric $g_{i}$ on $H^{3}(\mathbb{R})$ for which
$\|\text{v}_{1}\|_{g_{i}}^{4}=\|\text{v}_{2}\|_{g_{i}}^{4}=\|\text{v}_{3}\|_{g_{i}}=1/i^{4},\hskip 48.36958pt\langle\text{v}_{j_{1}},\text{v}_{j_{2}}\rangle_{g_{i}}=0\text{ for }j_{1}\neq j_{2},\hskip 14.22636pt$
and let $X_{i}$ be the quotient of $(H^{3}(\mathbb{R}),g_{i})$ by the action of $H^{3}(\mathbb{Z})$. Then
${\text{diam}}(X_{i})\to 0,\hskip 85.35826pt|\sec(X_{i})|\to 0.$
Note that ${\text{rank}}(\pi_{1}(X_{i}))={\text{rank}}(H^{3}(\mathbb{Z}))=3$. If one wants to construct finite-sheeted covering spaces $X_{i}^{\prime}\to X_{i}$ with
${\text{diam}}(X_{i}^{\prime})\leq D,\hskip 85.35826pt\mathcal{H}^{N}(X_{i}^{\prime})\geq\nu$
for some $D,\nu>0$, then one would need to unwind in all directions. That is, if $u_{j}:=\exp(\text{v}_{j})\in H^{3}(\mathbb{Z})$ for $j\in\{1,2,3\}$ and $X_{i}^{\prime}=(H^{3}(\mathbb{R}),g_{i})/\Gamma_{i}$ for some subgroup $\Gamma_{i}\leq H^{3}(\mathbb{Z})$, then there is a sequence $N_{i}\in\mathbb{N}$ going to infinity at least as fast as $i$ for which
$u_{1}^{n_{1}}u_{2}^{n_{2}}u_{3}^{n_{3}}\notin\Gamma_{i}$
for all $n_{1},n_{2},n_{3}\in\mathbb{Z}$ with $|n_{1}|,|n_{2}|\leq N_{i}$, $|n_{3}|\leq N_{i}^{4}$. The projection $\rho:H^{3}(\mathbb{Z})\to\mathbb{Z}^{2}$ given by
$u_{1}\mapsto(1,0),\hskip 42.67912ptu_{2}\mapsto(0,1),\hskip 42.67912ptu_{3}\mapsto(0,0)$
sends $\Gamma_{i}$ to a lattice $\Lambda_{i}$ generated by a basis $\{\text{w}_{1,i},\text{w}_{2,i}\}$. If we denote by $A_{i}$ the area of the parallelogram spanned by $\text{w}_{1,i}$ and $\text{w}_{2,i}$, then $A_{i}\geq N_{i}^{4}$. This follows from the fact that if $g_{j,i}\in\Gamma_{i}\cap\rho^{-1}(\text{w}_{j,i})$, then $g_{1,i},g_{2,i}]=u_{3}^{A_{i}}\in\Gamma_{i}$. From here, it is not hard to show that
${\text{diam}}(\mathbb{R}^{2}/\Lambda_{i})\geq\sqrt{A_{i}}/10\geq N_{i}^{2}/10.$
Finally, notice that the map $\overline{\rho}:H^{3}(\mathbb{R})\to\mathbb{R}^{2}$ given by
$\overline{\rho}\left(\begin{bmatrix}1&x&z\\ 0&1&y\\ 0&0&1\end{bmatrix}\right):=(x,y)$
induces an $i$-Lipschitz surjective map $X_{i}^{\prime}\to\mathbb{R}^{2}/\Lambda_{i}$. For $i$ large enough this implies
${\text{diam}}(X_{i}^{\prime})\geq{\text{diam}}(\mathbb{R}^{2}/\Lambda_{i})/i\geq N_{i}^{2}/10i\geq\sqrt{i}\to\infty,$
which is a contradiction.
[…]
The missing piece to prove Conjecture 2 is Lemma 4. That is, to prove the conjecture it would be enough to show that if in addition to the conditions of such lemma, one has
  • $X$ is a non-collapsed $\text{RCD}(K,N)$ space.
  • $Y$ is an almost flat nilpotent Lie group of dimension $N$.
  • $f$ satisfies the estimate
    $(1-\delta)\,d(x,y)^{1+\delta}\leq d(f(x),f(y))\leq(1+\delta)\,d(x,y)\,\text{ for all }x,y\in B_{10}(p).$
  • $G_{Y}$ acts freely on $Y$.
then the produced $\tilde{f}$ can be chosen to be locally bi-Hölder (and not just a homotopy equivalence). Arguing as in [44, Theorem 3.3] one can construct a global homeomorphism $\tilde{f}:X\to Y$ that is locally bi-Hölder, but making this function equivariant seems to be a very interesting technical challenge.
In principle, one would need to choose the bump functions $\rho_{g}$ more carefully, but this may not be enough as one may need to also modify the function $f$ as it is done in the smooth setting by Huang in [36]. Moreover, the exact same construction is possible as the groups we are dealing with are amenable, making the equivariant map constructed in [36, Theorems 1.3, 1.8] the perfect candidate to be the desired map.
2.1. Collapsed $\text{RCD}(K,N)$ spaces
For an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ with ${\text{diam}}(X)\leq\varepsilon(K,N)$, it follows from the proof of Theorem A that ${\text{rank}}(\pi_{1}(X))$ is at most the rectifiable dimension of $X$, which is an integer bounded above by $N$. One can ask whether the rigidity of Theorem A can be pushed to when ${\text{rank}}(\pi_{1}(X))$ equals the rectifiable dimension of $X$. This is not possible due to [77] (see also [39]).
[…]
For each $n\geq 3$, $\varepsilon>0$, there is an $\text{RCD}(-1,n+2)$ space $(X,\mathsf{d},\mathfrak{m})$ of rectifiable dimension equal to $n$ with
${\text{diam}}(X)\leq\varepsilon,\hskip 85.35826pt\pi_{1}(X)=\mathbb{Z}^{n},$
but no point of $X$ admits an open neighborhood homeomorphic to $\mathbb{R}^{n}$.
Proof.
This follows by performing the main construction of [77] to a flat torus of dimension $n$ and diameter $\leq\varepsilon$. ∎
It is however possible that if the dimension gap is small enough, the topological rigidity still holds.
[…]
For each $K\in\mathbb{R}$, $N\in\mathbb{N}$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N+\varepsilon)$ space with ${\text{diam}}(X)\leq\varepsilon$ and ${\text{rank}}(\pi_{1}(X))=N$, then $X$ is homeomorphic to an infranilmanifold.
2.2. Fibrations
Theorem A is a step towards the following conjectures which would extend to $\text{RCD}(K,N)$ spaces work of Naber–Zhang [55], Huang–Kong–Rong–Xu [37], and Huang–Wang [38].
[…]
Let $X$ be a geodesic space, $x\in X$, and $\varepsilon>0$. The fibered fundamental group $\Gamma_{\varepsilon}(x)$ is defined to be the image of the map $\pi_{1}(B_{\varepsilon}(x))\to\pi_{1}(B_{1}(x))$ induced by the inclusion.
[…]
There is $\varepsilon(K,N)>0$ such that the following holds. Let $X$ be a closed $m$-dimensional Riemannian manifold with
$|\sec(X)|\leq 1,\hskip 85.35826pt{\text{inj}}(X)\geq 1,$
and $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ a sequence of $\text{RCD}(K,N)$ spaces converging in the Gromov–Hausdorff sense to $X$. Assume that for all $i$ large enough and any $x\in X_{i}$ one has
${\text{rank}}(\Gamma_{\varepsilon}(x))=N-m.$
Then for large enough $i$ there are continuous Gromov–Hausdorff approximations $X_{i}\to X$ that are locally trivial fibrations with infranilmanifold fiber.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a closed $m$-dimensional Riemannian manifold $X$. Assume that for all $i$ one has
$\text{b}_{1}(X_{i})-\text{b}_{1}(X)=N-m.$
Then for large enough $i$ there are continuous Gromov–Hausdorff approximations $X_{i}\to X$ that are also locally trivial fibrations with fiber $\mathbb{T}^{N-m}$.
[…]
After the initial version of this paper appeared on arXiv, Wang [73] proved both Conjecture 2 and Conjecture 2.2 by building upon the techniques developed in this paper and in [76]. Moreover, while we pointed out in Proposition 2 that constructing non-collapsing finite-sheeted covers would not work in that setting, Wang found a way to overcome this difficulty and successfully implemented this technique to prove both conjectures.
3. Preliminary material
3.1. Notation
For a metric space $X$, we denote its group of isometries by $\text{Iso}(X)$. For $p\in X$, $r>0$, we denote the open ball of radius $r$ around $p$ by $B_{r}^{X}(p)$ or $B_{r}(p)$, depending on whether we want to emphasize the metric space used.
For $N\geq 0$, we denote by $\mathcal{H}^{N}$ the $N$-dimensional Hausdorff measure.
For $N\in\mathbb{N}$, we denote by $\mathbb{T}^{N}$ the torus
$\{(x_{1},y_{1},\ldots,x_{N},y_{N})\in\mathbb{R}^{2N}\,|\,x_{j}^{2}+y_{j}^{2}=1\text{ for each }j\}$
equipped with its standard flat metric and the $N$-dimensional Hausdorff measure.
$H_{m}(\cdot)$ denotes singular homology with integer coefficients. Singular simplices, singular chains, and the boundary operator $\partial$ are as defined in [33, Chapter 2].
For a vector space $V$ and a subset $S\subset V$, we denote by $\langle S\rangle\leq V$ the vector subspace generated by $S$. For $\text{v}\in V$, we also denote $\langle\{\text{v}\}\rangle$ by $\langle\text{v}\rangle$.
3.2. $\text{RCD}(K,N)$ spaces
Throughout this text, we assume the reader is familiar with the basic theory of $\text{RCD}(K,N)$ spaces, including the definition of Sobolev space $H^{1,2}$ and Laplace operator $\Delta$ on (infinitesimally Hilbertian) metric measure spaces. We refer the reader to [49, 65, 66, 2, 1] for the relevant definitions and to [19] for an overview of different equivalent definitions of $\text{RCD}(K,N)$ spaces.
We note that a large number of papers in the literature work with a condition known as $\text{RCD}^{\ast}(K,N)$, originally introduced in [5]. It is now known, thanks to the work of Cavalletti–Milman [11] and Li [48], that this condition is equivalent to the $\text{RCD}(K,N)$ condition.
While the topology of an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$ can be quite complicated (see for example [39, 77]), Wang showed that it is semi-locally-simply-connected [72]. Moreover, by the locality of the $\text{RCD}(K,N)$ condition [19, Section 3], lifting $\mathfrak{m}$ to the universal cover $\tilde{X}$ yields a measure $\tilde{\mathfrak{m}}$ that makes $\tilde{X}$ an $\text{RCD}(K,N)$ space (see [54]).
3.3. Equivariant Gromov–Hausdorff convergence
Throughout this paper, we assume the reader is familiar with the notions of Gromov–Hausdorff and pointed Gromov–Hausdorff convergence, as presented for example in [9].
We refer to [31] for the definition of convergence of maps and to [27] for the definition of equivariant Gromov–Hausdorff convergence. A feature of this frameword is that once one has pointed Gromov–Hausdorff convergence of proper spaces, compactness of groups of isometries comes for free [27, Proposition 3.6].
[…]
Let $(X_{i},p_{i})$ be a sequence of pointed proper metric spaces that converges in the pointed Gromov–Hausdorff sense to a pointed proper space $(X,p)$, and $\Gamma_{i}\leq\text{Iso}(X_{i})$ a sequence of groups of isometries. Then, after taking a subsequence, the triples $(X_{i},\Gamma_{i},p_{i})$ converge to a triple $(X,\Gamma,p)$ with $\Gamma\leq\text{Iso}(X)$.
Another feature of equivariant Gromov–Hausdorff convergence is its compatibility with quotients [25, Theorem 2.1].
[…]
Let $(X_{i},p_{i})$ be a sequence of pointed proper metric spaces and $\Gamma_{i}\leq\text{Iso}(X_{i})$ a sequence of groups of isometries. If the sequence of triples $(X_{i},\Gamma_{i},p_{i})$ converges to a triple $(X,\Gamma,p)$ in the equivariant Gromov–Hausdorff sense, then the seuquences of pointed quotient spaces $(X_{i}/\Gamma_{i},[p_{i}])$ converges in the pointed Gromov–Hausdorff sense to $(X/\Gamma,[p])$.
3.4. Measured Gromov–Hausdorff convergence
One of the main features of the class of $\text{RCD}(K,N)$ spaces is the measured Gromov–Hausdorff compactness. Throughout this paper, we use the extrinsic approach, introduced in [29] inspired on the original approach to pointed Gromov–Hausdorff convergence [31], and widely adopted in the literature (see for example [53, 3]). We say a triple $(X,\mathsf{d},\mathfrak{m})$ is a metric measure space if $(X,\mathsf{d})$ is a complete separable geodesic space and $\mathfrak{m}$ is a boundedly finite (finite on bounded sets) measure on $X$.
[…]
In a metric space $Z$, we say a sequence of boundedly finite measures $\mathfrak{m}_{i}$ converges weakly to a measure $\mathfrak{m}$ if for any bounded continuous function $f:Z\to\mathbb{R}$ with bounded support, one has
$\int_{Z}f\,\text{d}\mathfrak{m}_{i}\to\int_{Z}f\,\text{d}\mathfrak{m}.$
[…]
We say a sequence of compact metric measure spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ converges in the measured Gromov–Hausdorff sense to a compact metric measure space $(X,\mathsf{d},\mathfrak{m})$ if there is a separable metric space $(Z,\mathsf{d}_{Z})$ and isometric embeddings $\iota_{i}:X_{i}\to Z$, $\iota:X\to Z$ such that for every $\varepsilon>0$ there is $i_{0}(\varepsilon)\in\mathbb{N}$ such that for every $i\geq i_{0}$, $x\in X_{i}$, $y\in X$, there are $y^{\prime}\in X_{i}$ and $x^{\prime}\in X$ with
$\mathsf{d}_{Z}(\iota_{i}(x),\iota(x^{\prime})),\mathsf{d}_{Z}(\iota_{i}(y^{\prime}),\iota(y))<\varepsilon,$
and the sequence of measures $(\iota_{i})_{\ast}(\mathfrak{m}_{i})$ converges weakly to the measure $(\iota)_{\ast}(\mathfrak{m})$.
[…]
We say a sequence of pointed metric measure spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ converges in the pointed measured Gromov–Hausdorff sense to a pointed metric measure space $(X,\mathsf{d},\mathfrak{m},p)$ if there is a separable metric space $(Z,\mathsf{d}_{Z})$ and isometric embeddings $\iota_{i}:X_{i}\to Z$, $\iota:X\to Z$ such that for every $\varepsilon>0$ and $R>0$ there is $i_{0}(\varepsilon,R)\in\mathbb{N}$ such that for every $i\geq i_{0}$, $x\in B_{R}(p_{i})$, $y\in B_{R}(p)$, there are $y^{\prime}\in B_{R}(p_{i})$ and $x^{\prime}\in B_{R}(p)$ with
$\mathsf{d}_{Z}(\iota_{i}(x),\iota(x^{\prime})),\mathsf{d}_{Z}(\iota_{i}(y^{\prime}),\iota(y))<\varepsilon,$
and the sequence of measures $(\iota_{i})_{\ast}(\mathfrak{m}_{i})$ converges weakly to the measure $(\iota)_{\ast}(\mathfrak{m})$.
[…]
For each $i\in\mathbb{N}$, let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be an $\text{RCD}(K-\varepsilon_{i},N)$ space with $\varepsilon_{i}\to 0$. Then after normalizing the measures $\mathfrak{m}_{i}$, one can find a subsequence that converges in the pointed measured Gromov–Hausdorff sense to an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m},p)$.
Proof.
The fact that one can find a convergent subsequence follows from Gromov’s pre-compactness criterion [32, Proposition 5.2] combined with Prokhorov’s compactness theorem (see for example [68, Theorem 27.32]). The fact that the limit space is an $\text{RCD}(K,N)$ space was proven in [29] building upon [49, 65, 66, 2]. ∎
3.5. Dimension and volume convergence
Let $(X,\mathsf{d},\mathfrak{m})$ be an $\text{RCD}(K,N)$ space. In [8], Brué–Semola showed that there is a unique integer $n\leq N$ with the property that there is a set $\mathcal{R}_{n}\subset X$ of full measure such that for all $x\in\mathcal{R}_{n}$, the pointed Gromov–Hausdorff limit $\lim_{\lambda\to\infty}(\lambda X,x)$ exists and is isometric to $(\mathbb{R}^{n},0)$. The integer $n$ is called the rectifiable dimension of $X$.
[…]
Let $(X,\mathsf{d},\mathfrak{m})$ be an $\text{RCD}(K,N)$ space. Then the following are equivalent.
  1. (i)
    $X$ has rectifiable dimension $N$.
  2. (ii)
    $X$ has topological dimension $N$.
  3. (iii)
    $N\in\mathbb{N}$ and $X$ has Hausdorff dimension $N$.
  4. (iv)
    $N\in\mathbb{N}$ and $\mathfrak{m}=c\mathcal{H}^{N}$ for some $c>0$.
Proof.
Since reference measures on $\text{RCD}(K,N)$ spaces are non-zero and locally finite, (iv) implies (iii). By the work of De Philippis–Gigli [18, Theorem 1.12], (iv) implies (i). The fact that either (i) or (iii) implies (iv) was proven in [6, Theorem 2.20]. Since the topological dimension is always at most the Hausdorff dimension [40, Theorem VII 2], and the Hausdorff dimension of an $\text{RCD}(K,N)$ space is at most $N$, then (ii) implies (iii). The fact that (iv) implies (ii) follows from the work of Kapovitch–Mondino [44, Theorem 1.7]. ∎
An important feature of $\text{RCD}(K,N)$ spaces of maximal dimension is the continuity of volume, originally proven by Colding in the smooth setting [16].
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(K,N)$ spaces that converges in the pointed measured Gromov–Hausdorff sense to a pointed $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m},p)$ of rectifiable dimension $N$. Then
  1. (1)
    For $i$ large enough, $\mathfrak{m}_{i}$ is a constant multiple of $\mathcal{H}^{N}$.
  2. (2)
    $\lim_{i\to\infty}\mathcal{H}^{N}(B_{1}^{X_{i}}(p_{i}))=\mathcal{H}^{N}(B_{1}^{X}(p)).$
Proof.
The first claim follows from the lower semi-continuity of the rectifiable dimension [46, Theorem 1.5] combined with Theorem 3.5. The second claim follows from the work of De Philippis–Gigli on non-collapsed $\text{RCD}(K,N)$ spaces [18, Theorem 1.3]. ∎
A consequence of Theorem 3.5 is that non-collapsing sequences of $\text{RCD}(K,N)$ spaces admit no small groups of isometries.
[…]
Let $(X_{i},p_{i})$ be a sequence of proper geodesic spaces. We say a sequence of groups $W_{i}\leq\text{Iso}(X_{i})$ consists of small subgroups if for each $R>0$ we have
$\lim_{i\to\infty}\,\sup_{g\in W_{i}}\,\sup_{x\in B_{R}(p_{i})}d(gx,x)=0.$
Equivalently, the groups $W_{i}$ are small if they converge to the trivial group in the equivariant Gromov–Hausdorff sense and
$\lim_{i\to\infty}\sup_{g\in W_{i}}d(gp_{i},p_{i})<\infty.$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(K,N)$ spaces that converges in the pointed measured Gromov–Hausdorff sense to a pointed $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m},p)$ of rectifiable dimension $N$. Then any sequence of small subgroups $W_{i}\leq\text{Iso}(X_{i})$ is eventually trivial.
Proof.
Working by contradiction, after passing to a subsequence, we assume $W_{i}$ is non-trivial for each $i$. For $i$ large enough, the closure $\overline{W_{i}}$ is a non-trivial compact Lie group, so it contains a non-trivial finite group $F_{i}\leq\overline{W_{i}}$. By Theorem 3.5(1), the groups $\text{Iso}(X_{i})$ are measure preserving, so by the work of Galaz–Kell–Mondino–Sosa [28, Theorem 1.1] the pushforward measure $\overline{\mathfrak{m}}_{i}$ makes $X_{i}/F_{i}$ an $\text{RCD}(K,N)$ space.
Since the groups $F_{i}$ are small, the sequence $(X_{i}/F_{i},d_{X_{i}/F_{i}},\overline{\mathfrak{m}}_{i},[p_{i}])$ also converges in the pointed measured Gromov–Hausdorff sense to $(X,\mathsf{d},\mathfrak{m},p)$. However, applying Theorem 3.5(2) twice we get
$\mathcal{H}^{N}(B_{1}^{X}(p))=\lim_{i\to\infty}\mathcal{H}^{N}(B_{1}^{X_{i}/F_{i}}([p_{i}]))=\lim_{i\to\infty}\frac{1}{|F_{i}|}\mathcal{H}^{N}(B_{1}^{X_{i}}(p_{i}))\leq\frac{1}{2}\mathcal{H}^{N}(B_{1}^{X}(p)),$
which is a contradiction. ∎
3.6. Topological stability
The following is a special case of [44, Theorem 2.9].
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. Then there is a sequence of pointed Gromov–Hausdorff approximations $f_{i}:(X_{i},p_{i})\to(\mathbb{R}^{N},0)$ such that for all $R>\delta>0$, the restriction of $f_{i}$ to $B_{R}(p_{i})$ is a homeomorphism with its image for $i$ large enough depending on $R,\delta$, and
$B_{R-\delta}(0)\subset f_{i}(B_{R}(p_{i})).$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. If there are groups $\Gamma_{i}\leq\text{Iso}(X_{i})$ with
$\limsup_{i\to\infty}\,{\text{diam}}(X_{i}/\Gamma_{i})<\infty,$
then $X_{i}$ is a topological manifold of dimension $N$ for $i$ large enough.
Proof.
Let $R>\limsup_{i}{\text{diam}}(X_{i}/\Gamma_{i})$. By Theorem 3.6, for $i$ large enough each point in $B_{R}(p_{i})$ has a neighborhood homeomorphic to $\mathbb{R}^{N}$. Since also for $i$ large enough and each $x\in X_{i}$ there is an isometry $g\in\Gamma_{i}$ with $gx\in B_{R}(p_{i})$ the result follows. ∎
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ be a sequence of pointed $\text{RCD}(-\frac{1}{i},N)$ spaces such that the sequence $(X_{i},p_{i})$ converges in the pointed Gromov–Hausdorff sense to $(\mathbb{R}^{N},0)$. Then for any $\varepsilon$, the following holds for $i$ large enough:
  • For all $r\in(0,1]$, $x\in B_{1}(p_{i})$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
Proof.
Let $f_{i}:X_{i}\to X$ be the Gromov–Hausdorff approximations given by Theorem 3.6. Assuming the result fails, after passing to a subsequence, there are sequences $x_{i}\in B_{1}(p_{i})$, $r_{i}\in(0,1]$ for which the ball $B_{r_{i}}(x_{i})$ is not contractible in $B_{r_{i}+\varepsilon}(x_{i})$. After passing further to a subsequence, we can assume $r_{i}\to r$ for some $r\in[0,1]$ and $f_{i}(x_{i})\to x$ for some $x\in B_{1}(0)$.
Then for $i$ large enough, $B_{r+\varepsilon/4}(x_{i})$ is not contractible in $B_{r+3\varepsilon/4}(x_{i})$. However, this is false, as for $i$ large enough one has
$f_{i}(B_{r+\varepsilon/4}(x_{i}))\subset B_{r+\varepsilon/2}(x)\subset f_{i}(B_{r+3\varepsilon/4}(x_{i})),$
so we can pull back via $f_{i}$ a nullhomotopy of $B_{r+\varepsilon/2}(x)$. ∎
3.7. Analysis and Spectral Estimates
Let $(X,\mathsf{d},\mathfrak{m})$ be a compact $\text{RCD}(K,N)$ space of diameter $\leq D$. The spectrum of its Laplacian $\Delta$ is discrete and non-positive, and one has an orthogonal direct sum decomposition
$H^{1,2}(X)=\bigoplus_{\lambda\geq 0}\{f\in H^{1,2}(X)|\Delta f=-\lambda f\}.$
For an eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$, one has
$\|f\|_{H^{1,2}}^{2}=(1+\lambda)\|f\|_{L^{2}}^{2}.$
Hence by [4, Proposition 7.1] and the discussion preceding it, there is $C(K,N,D)>0$ for which
(3.1) $\hskip 17.07182pt\|f\|_{L^{\infty}}\leq C\lambda^{(N-2)/4}\|f\|_{H^{1,2}},\hskip 34.14322pt\|\nabla f\|_{L^{\infty}}\leq C\lambda^{N/4}\|f\|_{H^{1,2}}.$
The Laplacian spectrum is continuous with respect to measured Gromov–Hausdorff convergence. This continuity was originally conjectured for Riemannian manifolds by Fukaya [26], proven in that context and more general ones by Cheeger–Colding [13], and extended to $\text{RCD}(K,N)$ spaces by Gigli–Mondino–Savare [29] (see also [3] for an extension in non-compact spaces).
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of $\text{RCD}(K,N)$ spaces converging in the measured Gromov–Hausdorff sense to an $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$, and let $(Z,\mathsf{d}_{Z})$, $\iota_{i}:X_{i}\to Z$, $\iota:X\to Z$ be as in Definition 3.4. Given a sequence $f_{i}\in H^{1,2}(X_{i})$, we say it converges in $H^{1,2}$ to $f\in H^{1,2}(X)$ if the sequence of measures $(\iota_{i})_{\ast}(f_{i}\mathfrak{m}_{i})$ converges weakly to $(\iota)_{\ast}(f\mathfrak{m})$ and
$\lim_{i\to\infty}\|f_{i}\|_{H^{1,2}}=\|f\|_{H^{1,2}}.$
We say $f_{i}$ converges uniformly to $f$ if for any choice of points $x_{i}\in X_{i}$, $x\in X$ with $\iota_{i}(x_{i})\to\iota(x)$, one has $f_{i}(x_{i})\to f(x)$.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Then
  1. (1)
    For each sequence of eigenfunctions $f_{i}\in H^{1,2}(X_{i})$ with
    $\Delta f_{i}=-\lambda_{i}f_{i},\hskip 36.98866pt\limsup_{i\to\infty}\lambda_{i}<\infty,\hskip 28.45274pt\limsup_{i\to\infty}\|f_{i}\|_{H^{1,2}}<\infty,$
    there is an eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$ such that after passing to a subsequence, $\lambda_{i}\to\lambda$ and $f_{i}$ converges to $f$ both uniformly and in $H^{1,2}$.
  2. (2)
    For each eigenfunction $f\in H^{1,2}(X)$ with $\Delta f=-\lambda f$, there is a sequence of eigenfunctions $f_{i}\in H^{1,2}(X_{i})$ with $\Delta f_{i}=-\lambda_{i}f_{i}$ such that $\lambda_{i}\to\lambda$ and $f_{i}$ converges to $f$ both uniformly and in $H^{1,2}$.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Then the sequence of Laplace spectra of the spaces $X_{i}$ counted with multiplicities as measures on $\mathbb{R}$ converge weakly to the Laplace spectrum of $X$.
3.8. Canonical homeomorphisms
Eigenfunctions provide means of constructing bi-Hölder homeomorphisms between suitable non-collapsed $\text{RCD}(K,N)$ spaces. Theorem 3.8 below was obtained in [35, Theorem 1.2.2], inspired by the work of Wang–Zhao [70], and built upon the tools developed by Cheeger–Jiang–Naber [15].
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces. A sequence of maps $\Phi_{i}:X_{i}\to\mathbb{R}^{k}$ is said to be equi-regular if for any smooth function $f\in C^{\infty}(\mathbb{R}^{k})$, the composition $f\circ\Phi_{i}:X_{i}\to\mathbb{R}$ is in the domain of the Laplacian for all $i$ and
$\sup_{i}\|\Delta(f\circ\Phi_{i})\|_{\infty}<\infty.$
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces that converges in the measured Gromov–Hausdorff sense to a compact $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m})$. Assume $(X,\mathsf{d})$ is isometric to a smooth $N$-dimensional Riemannian manifold, and consider a smooth isometric embedding $\Phi:X\to\mathbb{R}^{k}$, together with an open neighborhood $U\subset\mathbb{R}^{k}$ and a $C^{1,1}$-Riemannian submersion $\pi:U\to\Phi(X)$ with $\pi|_{\Phi(X)}=\text{Id}_{\Phi(X)}$. Then if $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$ is a sequence of equi-regular maps converging uniformly to $\Phi$ such that the family of functions $\Delta\varphi_{j,i}$ is equi-Lipschitz, then the sequence of maps
$F_{i}:=(\Phi^{-1}\circ\pi\circ\Phi_{i}):X_{i}\to X$
satisfies
$(1-\varepsilon_{i})\,\mathsf{d}_{i}(x,y)^{1+\varepsilon_{i}}\leq\mathsf{d}(F_{i}(x),F_{i}(y))\leq(1+\varepsilon_{i})\,\mathsf{d}_{i}(x,y)\,\text{ for all }x,y\in X_{i}$
for some $\varepsilon_{i}\to 0$. Furthermore, if the spaces $X_{i}$ are smooth Riemannian manifolds, and $\Phi_{i}$ is smooth for each $i$, then $F_{i}$ is a diffeomorphism for $i$ large enough.
[…]
Let $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ be a sequence of compact $\text{RCD}(K,N)$ spaces of uniformly bounded diameter and $\varphi_{j,i}\in H^{1,2}(X_{i})$ a family of eigenfunctions with $\Delta\varphi_{j,i}=-\lambda_{j,i}\varphi_{j,i}$ for $j\in\{1,\ldots,k\}$ and assume
$\sup_{i,j}\,\lambda_{j,i}<\infty,\hskip 85.35826pt\sup_{i,j}\,\|\varphi_{j,i}\|_{2}<\infty.$
Consider the map $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$. Since for any smooth function $f\in C^{\infty}(\mathbb{R}^{k})$, one has
$\Delta(f\circ\Phi_{i})=\sum_{j=1}^{k}\left[\frac{\partial f}{\partial x_{j}}\circ\Phi_{i}\right]\Delta\varphi_{j,i}+\sum_{j,\ell=1}^{k}\left[\frac{\partial^{2}f}{\partial x_{j}\partial x_{\ell}}\circ\Phi_{i}\right]\langle\nabla\varphi_{j,i},\nabla\varphi_{\ell,i}\rangle,$
then by (3.1), the maps $\Phi_{i}=(\varphi_{1,i},\ldots,\varphi_{k,i}):X_{i}\to\mathbb{R}^{k}$ are equi-regular and equi-Lipschitz.
3.9. Group theory
Recall that a finitely generated group $\Lambda$ is said to be polycyclic if there is a sequence of subgroups
$1=\Lambda_{0}\trianglelefteq\cdots\trianglelefteq\Lambda_{m}=\Lambda$
with $\Lambda_{j}/\Lambda_{j-1}$ cyclic for each $j$. In such a case, ${\text{rank}}(\Lambda)$ is defined to be the number of $j$’s for which the quotient $\Lambda_{j}/\Lambda_{j-1}$ is isomorphic to $\mathbb{Z}$. By Schreier Theorem [23] this number is well defined. The following properties come directly from the definition.
[…]
If $\Lambda^{\prime}\leq\Lambda$ is a finite index subgroup then ${\text{rank}}(\Lambda^{\prime})={\text{rank}}(\Lambda)$.
[…]
For a short exact sequence
$1\to K\to\Lambda\to\Gamma\to 1$
with $\Lambda$ a polycyclic group, we have
${\text{rank}}(\Lambda)={\text{rank}}(K)+{\text{rank}}(\Gamma).$
For an arbitrary finitely generated group $G$, ${\text{rank}}(G)$ is defined as the infimum of ${\text{rank}}(\Lambda)$ among finite index polycyclic subgroups $\Lambda\leq G$. By Proposition 3.9 if $G$ is polycyclic there is no conflict between the distinct definitions of ${\text{rank}}(G)$.
[…]
Let $G$ be a group, $\Gamma$ a polycyclic group, and $\theta:G\to\Gamma$ a surjective morphism with finite kernel. Then ${\text{rank}}(G)={\text{rank}}(\Gamma)$.
Proof.
Let $\Lambda\leq G$ be the kernel of the adjoint action on ${\text{Ker}}(\theta)\trianglelefteq G$. Then there is a short exact sequence
$1\to{\text{Ker}}(\theta)\cap\Lambda\to\Lambda\to\theta(\Lambda)\to 1.$
${\text{Ker}}(\theta)\cap\Lambda$ is finite abelian, hence polycyclic. Since ${\text{Ker}}(\theta)\cap\Lambda$ and $\theta(\Lambda)$ are polycyclic, so is $\Lambda$. Then by Propositions 3.9 and 3.9;
${\text{rank}}(G)={\text{rank}}(\Lambda)={\text{rank}}(\theta(\Lambda))={\text{rank}}(\Gamma).$
The ensuing lemma follows from classic simultaneous diagonalization.
[…]
Let $G$ be an abelian group acting by linear isometries on a finite dimensional real vector space $V$ of dimension $2N$. Assume $V=\oplus_{\ell=1}^{m}E_{\ell}$ for some $G$-invariant subspaces $E_{\ell}$. Then $V$ admits a (not necessarily unique!) decomposition $V=\oplus_{j=1}^{N}W_{j}$ with each $W_{j}$ 2-dimensional, $G$-invariant, and satisfying
$W_{j}=\bigoplus_{\ell=1}^{m}\,(E_{\ell}\cap W_{j}).$
[…]
For a group $G$ and two subgroups $H_{1},H_{2}\leq G$, we denote by $H_{1},H_{2}]\leq G$ the subgroup generated by elements of the form $h_{1}h_{2}h_{1}^{-1}h_{2}^{-1}$ with $h_{1}\in H_{1}$, $h_{2}\in H_{2}$. We set $G^{(0)}:=G$, and define inductively $^{(j+1)}:=[G^{(j)},G]. GG$ is said to be nilpotent if $G^{(s)}$ is the trivial group for some $s\in\mathbb{N}$.
3.10. Limits of almost homogeneous spaces
In a portion of this paper, we follow the construction performed in [76], which is an adaptation to metric spaces of a construction by Breuillard–Green–Tao in [7]. This construction begins with the following result, which establishes that discrete groups that act almost transitively on proper geodesic spaces are virtually nilpotent.
[…]
Let $(X_{i},p_{i})$ be a sequence of pointed proper geodesic spaces that converges to a pointed proper semi-locally-simply-connected geodesic space $(X,p)$ in the pointed Gromov–Hausdorff sense, and $G_{i}\leq\text{Iso}(X_{i})$ a sequence of discrete groups with ${\text{diam}}(X_{i}/G_{i})\to 0$. Then there exists $s\in\mathbb{N}$ and a sequence of finite-index subgroups $G_{i}^{\prime}\leq G_{i}$ with
(3.2) $\lim_{i\to\infty}\,\,\,\sup_{x\in X_{i}}\,\sup_{g\in(G_{i}^{\prime})^{(s)}}d(gx,x)=0,$
(3.3) $\limsup_{i\to\infty}\,[G_{i}:G_{i}^{\prime}]<\infty.$
Proof.
By [76, Theorem 4.1], $X$ is a nilpotent Lie group equipped with a sub-Finsler metric. Sub-Finsler metrics are locally doubling [56], so [76, Remark 3.4] applies and [76, Lemma 3.2] provides the desired groups $G_{i}^{\prime}$. ∎
It is well known that if a Riemannian manifold admits a nilpotent group of isometries acting transitively, then such group acts freely [74]. Conditions (3.2) and (3.3) imply that the groups $G_{i}^{\prime}$ are almost nilpotent and act almost transitively, so it is not surprising that they act almost freely. This is the content of the following lemma.
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$, be as in Lemma 3.10. Then after passing to a subsequence the groups $G_{i}^{\prime}$ converge in the equivariant Gromov–Hausdorff sense to a connected nilpotent group $\Gamma\leq\text{Iso}(X)$ acting freely and transitively.
Proof.
By Theorem 3.3, after passing to a subsequence the groups $G_{i}^{\prime}$ converge to a group $\Gamma\leq\text{Iso}(X)$. By the proof of [76, Theorem 3.1], there is a nilpotent subgroup $\mathcal{O}\leq\Gamma$ acting freely and transitively on $X$, so we can identify $\mathcal{O}$ with $X$. By [76, Theorem 4.1], $\mathcal{O}$ is a Lie group, and by [76, Proposition 5.2], $\Gamma$ acts freely on $X$. Since the action of $\mathcal{O}$ is transitive, the only possibility is that $\Gamma=\mathcal{O}$. ∎
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$, be as in Lemma 3.10. If the groups $G_{i}$ are such that any sequence of small subgroups is eventually trivial, then $G_{i}^{\prime}$ acts freely for $i$ large enough.
Proof.
Assuming the contrary, after passing to a subsequence there are $x_{i}\in X_{i}$ with non-trivial stabilizers $K_{i}\leq G_{i}^{\prime}$. Notice that since the groups $G_{i}^{\prime}$ act almost transitively, we can assume $x_{i}\in B_{1}(p_{i})$ for $i$ large enough. After again passing to a subsequence, $x_{i}$ converges to a point $x\in X$ and the groups $K_{i}$ converge to a group $K\leq\Gamma$ contained in the stabilizer of $x$. By Lemma 3.10, $K$ is trivial, so the groups $K_{i}$ are small, contradicting the hypothesis. ∎
Lemma 3.10 is particularly powerful when paired with Lemma 3.10, which follows from the proof of [67, Corollary 4.2.7].
[…]
Let $\Gamma\leq\text{Iso}(\mathbb{R}^{n})$ be a connected nilpotent group of isometries acting transitively. Then $\Gamma$ is the group of translations in $\mathbb{R}^{n}$.
Lemma 3.10 below gives an explicit description of the groups $G_{i}^{\prime}$.
[…]
Let $G$ be a group, $u_{1},u_{2},\ldots,u_{r}\in G$, and $N_{1},N_{2},\ldots,N_{r}$ $\in\mathbb{R}^{+}$. The set $P(u_{1},\ldots,u_{r};N_{1},\ldots,N_{r})\subset G$ is defined to be the set of elements that can be expressed as words in the $u_{i}$’s and their inverses such that the number of appearances of $u_{i}$ and $u_{i}^{-1}$ is not more than $N_{i}$. We then say that $P(u_{1},\ldots,u_{r};N_{1},\ldots,N_{r})$ is a nilprogression in $C$-normal form for some $C>0$ if it also satisfies the following properties:
  1. (1)
    For all $1\leq i\leq j\leq r$, and all choices of signs, we have
    $u_{i}^{\pm 1},u_{j}^{\pm 1}]\in P\left(u_{j+1},\ldots,u_{r};\dfrac{CN_{j+1}}{N_{i}N_{j}},\ldots,\dfrac{CN_{r}}{N_{i}N_{j}}\right).$
  2. (2)
    The expressions $u_{1}^{n_{1}}\ldots u_{r}^{n_{r}}$ represent distinct elements as $n_{1},\ldots,n_{r}$ range over the integers with $|n_{1}|\leq N_{1}/C,\ldots,|n_{r}|\leq N_{r}/C$.
  3. (3)
    One has
    $\frac{1}{C}(2\lfloor N_{1}\rfloor+1)\cdots(2\lfloor N_{r}\rfloor+1)\leq|P|\leq C(2\lfloor N_{1}\rfloor+1)\cdots(2\lfloor N_{r}\rfloor+1).$
For a nilprogression $P$ in $C$-normal form, and $\varepsilon\in(0,1)$, the set $P(u_{1},\ldots,u_{r};$ $\varepsilon N_{1},\ldots,\varepsilon N_{r})$ also satisfies conditions (1) and (2), and we denote it by $\varepsilon P$. We define the thickness of $P$ as the minimum of $N_{1},\ldots,N_{r}$ and we denote it by $\text{thick}(P)$. The set $\{u_{1}^{n_{1}}\ldots u_{r}^{n_{r}}||n_{i}|\leq N_{i}/C\}$ is called the grid part of $P$, and is denoted by $G(P)$.
[…]
Let $(X_{i},p_{i})$, $(X,p)$, $G_{i}$, $G_{i}^{\prime}$ be as in Lemma 3.10 and $n$ the topological dimension of $X$. Then for $i$ large enough, there are torsion-free nilpotent groups $\tilde{\Gamma}_{i}$ generated by elements $\tilde{u}_{1,i},\ldots,\tilde{u}_{n,i}\in\tilde{\Gamma}_{i}$, and $N_{1,i},\ldots,N_{n,i}\in\mathbb{R}^{+}$ with the following properties:
  1. (i)
    The set map $\alpha_{i}:\mathbb{Z}^{n}\to\tilde{\Gamma}_{i}$ given by
    $\alpha_{i}(m_{1},\ldots,m_{n})=\tilde{u}_{1,i}^{m_{1}}\cdots\tilde{u}_{n,i}^{m_{n}}$
    is a bijection. Moreover, there are polynomials $Q_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ of degree $\leq d(n)$ such that the product in $\tilde{\Gamma}_{i}$ is given by
    $g_{1}\cdot g_{2}=\alpha_{i}(Q_{i}(\alpha_{i}^{-1}(g_{1}),\alpha_{i}^{-1}(g_{2}))).$
  2. (ii)
    There are small normal subgroups $W_{i}\trianglelefteq G_{i}^{\prime}$ and surjective group morphisms
    $\Phi_{i}:\tilde{\Gamma}_{i}\to\Gamma_{i}:=G_{i}^{\prime}/W_{i}$
    with ${\text{Ker}}(\Phi_{i})$ containing an isomorphic copy of $\pi_{1}(X)$.
  3. (iii)
    There is $C>0$ such that if $u_{j,i}:=\Phi_{i}(\tilde{u}_{j,i})$ for each $j\in\{1,\ldots,n\}$, the set
    $P_{i}:=P(u_{1,i},\ldots,u_{n,i};N_{1,i},\ldots,N_{n,i})\subset\Gamma_{i}$
    is a nilprogression in $C$-normal form with $\text{thick}(P_{i})\to\infty$.
  4. (iv)
    For each $\varepsilon>0$ there is $\delta>0$ such that
    $G(\delta P_{i})\subset\{g\in\Gamma_{i}\,|\,$ $d(g[p_{i}],[p_{i}])\leq\varepsilon\},$
    $\{g\in\Gamma_{i}\,|\,d(g[p_{i}],[p_{i}$ $])\leq\delta\}\subset G(\varepsilon P_{i})$
    for $i$ large enough, where we are considering the action of $\Gamma_{i}$ on $X_{i}/W_{i}$,
Proof.
In [76, Section 6] first the small subgroups $W_{i}\trianglelefteq G_{i}^{\prime}$ are constructed (using the escape norm from [7]) in such a way that the quotients $G_{i}^{\prime}/W_{i}$ have no small subgroups. Then in [7, Section 9] the nilprogressions $P_{i}\subset\Gamma_{i}$ are obtained via a short basis construction. (iii) and the first item of (iv) follow from [7, Theorem 9.3] and its proof, while the second item of (iv) is proven in [76, Theorem 7.1]. In [76, Section 8], (i) is obtained by an application of Malcev Theorem, and (ii) follows from [76, Proposition 9.1]. Note that although the results in [7, 76] are stated in terms of ultralimits and non-standard analysis, they are equivalent to the above formulations by Łos’s Theorem (see for example [7, Appendix C]). ∎
[…]
Combining Lemma 3.10 with Lemma 3.9 and Proposition 3.9 one obtains
(3.4) $\begin{split}{\text{rank}}(G_{i}^{\prime})&={\text{rank}}(\Gamma_{i})\\ &={\text{rank}}(\tilde{\Gamma}_{i})-{\text{rank}}({\text{Ker}}(\Phi_{i}))\\ &\leq n.\end{split}$
If equality holds throughout (3.4), then ${\text{rank}}({\text{Ker}}(\Phi_{i}))=0$, and since the groups $\tilde{\Gamma}_{i}$ are torsion-free, that would imply ${\text{Ker}}(\Phi_{i})=0$, so $X$ is simply connected and $\tilde{\Gamma}_{i}=\Gamma_{i}$.
[…]
If the groups $G_{i}^{\prime}$ admit no small subgroups, then $\Gamma_{i}=G_{i}^{\prime}$ for $i$ large enough.
4. Topological Lemmas
In general, when two pointed proper metric spaces $(X,p)$, $(Y,q)$ are pointed Gromov–Hausdorff close, there is no global approximation $X\to Y$. However, when a group $G$ acts co-compactly on both spaces and the triples $(X,p,G)$, $(Y,q,G)$ are equivariantly Gromov–Hausdorff close, it is possible to construct a global almost $G$-equivariant approximation $X\to Y$ (see for example [71, Theorem 4.1]).
[…]
For each $\varepsilon>0$, there is $\delta>0$ such that the following holds. Let $X$, $Y$ be proper geodesic spaces, $G_{X}\leq\text{Iso}(X)$, $G_{Y}\leq\text{Iso}(Y)$ discrete groups, $p\in X$, $q\in Y$, a group isomorphism $\varphi:G_{X}\to G_{Y}$, and a map $f:B_{10}(p)\to B_{10}(q)$ such that $f(p)=q$ and
  1. (1)
    ${\text{diam}}(X/G_{X}),$ ${\text{diam}}(Y/G_{Y})<1$.
  2. (2)
    For all $x_{1},x_{2}\in B_{10}(p)$,
    $|d(f(x_{1}),f(x_{2}))-d(x_{1},x_{2})|\leq\delta.$
  3. (3)
    For all $y\in B_{1}(q)$, there is $x\in B_{2}(p)$ with
    $d(f(x),y)\leq\delta.$
  4. (4)
    If $x\in B_{10}(p)$, $h\in G_{Y}$ are such that $hf(x)\in B_{5}(q)$, then
    $\varphi^{-1}(h)x\in B_{10}(p).$
  5. (5)
    If $x\in B_{10}(p)$, $g\in G_{X}$ are such that $gx\in B_{10}(p)$, then
    $d(f(gx),\varphi(g)f(x))\leq\delta.$
Then there is $\tilde{f}:X\to Y$ such that
  1. (i)
    For all $x\in X$, $g\in G_{X}$,
    $d(\varphi(g)\tilde{f}(x),\tilde{f}(gx))\leq\varepsilon.$
  2. (ii)
    For all $x_{1},x_{2}\in X$ with $\min\{d(x_{1},x_{2}),d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))\}\leq 2$ one has
    $|d(\tilde{f}(x_{1})\tilde{f}(x_{2}))-d(x_{1},x_{2})|\leq\varepsilon.$
  3. (iii)
    For all $y\in Y$, there is $x\in X$ with $d(\tilde{f}x,y)\leq\varepsilon$.
Moreover, if
(4.1) $Y\text{ is Riemannian,}\hskip 56.9055pt|\sec(Y)|\leq 1,\hskip 56.9055pt{\text{inj}}(Y)\geq 1,$
and $f$ is continuous, then $\tilde{f}$ can be taken to be continuous and equivariant; that is, for all $x\in X$, $g\in G_{X}$,
$\varphi(g)\tilde{f}(x)=\tilde{f}(gx).$
Proof.
For each $g\in G_{X}$, set $\theta_{g}:X\to\mathbb{R}$ as
$\theta_{g}(x):=\max\{0,2-d(x,gp)\}.$
Then define $\Theta:X\to\mathbb{R}$ as
$\Theta(x)=\sum_{g\in G_{X}}\theta_{g}(x),$
which by (1) is always positive. Also for each $g\in G_{X}$ define $\rho_{g}:X\to\mathbb{R}$ as
$\rho_{g}(x):=\theta_{g}(x)/\Theta(x),$
and with this, the map $\mu:X\to\text{Prob}(Y)$ as
$\mu_{x}:=\sum_{g\in G_{X}}\rho_{g}(x)[[\varphi(g)(f(g^{-1}x))]],$
where for each $y\in Y$, we denote by $y]]\in\text{Prob}(Y)$ the delta measure supported at $y$. Notice that $\rho_{g}(x)\neq 0$ means that $g^{-1}x\in B_{1}(p)$, so $\mu_{x}$ is well defined even if $f(g^{-1}x)$ is not for most $g\in G_{X}$. It follows from the definitions that
$\rho_{g}(hx)=\rho_{h^{-1}g}(x)\text{ for all }x\in X,\,g,h\in G_{X}.$
Then
(4.2) $\begin{split}\mu_{hx}&=\sum_{g\in G_{X}}\rho_{g}(hx)[[\varphi(g)(f(g^{-1}hx))]]\\ &=\sum_{g\in G_{X}}\rho_{h^{-1}g}(x)[[\varphi(g)(f((h^{-1}g)^{-1}x))]]\\ &=\sum_{u\in G_{X}}\rho_{u}(x)[[\varphi(hu)(f(u^{-1}x))]]\\ &=\varphi(h)_{\ast}\sum_{u\in G_{X}}\rho_{u}(x)[[\varphi(u)(f(u^{-1}x))]]\\ &=\varphi(h)_{\ast}\mu_{x},\end{split}$
so $\mu$ is an equivariant map. Also, if $x\in X$, $g,h\in G_{X}$ are such that $\rho_{g}(x),\rho_{h}(x)>0$, then $g^{-1}(x),h^{-1}(x)\in B_{2}(p)$, and by (5) we have
$d(\varphi(g)f($ $g^{-1}x),\varphi(h)f(h^{-1}x))=d(f(g^{-1}x),\varphi(g^{-1}h)f(h^{-1}x))$
$=d(f((g^{-1}h)h^{-1}x),\varphi(g^{-1}h)f(h^{-1}x))$
$\leq\delta,$
so
(4.3) ${\text{diam}}({\text{supp}}(\mu_{x}))\leq\delta\text{ for all }x\in X.$
Then define $\tilde{f}:X\to Y$ as any map with
(4.4) $\tilde{f}(x)\in{\text{supp}}(\mu_{x}).$
By (4.3), different choices for $\tilde{f}$ yield functions that are uniformly $\delta$-close, and (i) follows from (4.2). Now if $x_{1},x_{2}\in X$ are such that $d(x_{1},x_{2})\leq 5,$ take $g_{1},g_{2}\in G_{X}$ with $g^{-1}_{1}x_{1},g^{-1}_{2}x_{2}\in B_{1}(p)$. Then $g_{1}^{-1}x_{2}\in B_{10}(p)$ so by (2) and (5) we have
$d(\varphi(g_{1})f($ $g_{1}^{-1}x_{1}),\varphi(g_{2})f(g_{2}^{-1}x_{2}))=d(f(g_{1}^{-1}x_{1}),\varphi(g_{1}^{-1}g_{2})f(g_{2}^{-1}x_{2}))$
$=d(f(g_{1}^{-1}x_{1}),f(g_{1}^{-1}x_{2}))\pm\delta$
$=d(g_{1}^{-1}x_{1},g_{1}^{-1}x_{2})\pm 2\delta$
$=d(x_{1},x_{2})\pm 2\delta.$
Combining this with (4.3) we obtain
(4.5) $|d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))-d(x_{1},x_{2})|\leq 4\delta.$
On the other hand, if $x_{1},x_{2}$ are such that $d(\tilde{f}(x_{1}),\tilde{f}(x_{2}))\leq 2$, then there are $g_{1},g_{2}\in G_{X}$ such that $\rho_{g_{1}}(x_{1}),\rho_{g_{2}}(x_{2})>0$, and
$d(\varphi(g_{1})f(g_{1}^{-1}x_{1}),\varphi(g_{2})f(g_{2}^{-1}x_{2}))\leq 3.$
Using (2) and the fact that $f(p)=q$, this implies
$d(\varphi(g_{2}^{-1}g_{1})f(g_{1}^{-1}x_{1}),q)\leq 5,$
so by (4), we get $g_{2}^{-1}g_{1}g_{1}^{-1}x_{1}=g_{2}^{-1}x_{1}\in B_{10}(p)$. Hence by (2),
$d(x_{1},x_{2})$ $=d(g_{2}^{-1}x_{1},g_{2}^{-1}x_{2})$
$\leq d(f(g_{2}^{-1}x_{1}),f(g_{2}^{-1}x_{2}))+\delta$
$\leq d(\varphi(g_{2}^{-1}g_{1})f(g_{1}^{-1}x_{1}),f(g_{2}^{-1}x_{2}))+2\delta$
$\leq 5,$
so (4.5) applies, proving (ii). For $y\in Y$, by (1) there is $g\in G_{X}$ with
$\varphi(g)^{-1}y\in B_{1}(q).$
By (3), there is $x\in B_{2}(p)$ with $d(f(x),\varphi(g)^{-1}y)\leq\delta$. Since $\rho_{g}(gx)>0$, we have
$d(\tilde{f}(gx),y)\leq d(\varphi(g)f(g^{-1}gx),y)+\delta\leq 2\delta,$
establishing (iii).
Now assume $Y$ satisfies (4.1) and $f$ is continuous. Then instead of defining $\tilde{f}$ by (4.4), define $\tilde{f}(x)\in Y$ to be the center of mass of $\mu_{x}$. As the functions $\rho_{g}$ are all continuous, $\mu_{x}$ depends continuously and equivariantly on $x$, so the result follows. ∎
When two geodesic spaces $X$ and $Y$ are Gromov–Hausdorff close, very little can be said about their topologies (see for example [21]). However, if $X$ satisfies nice local contractibility properties and admits a continuous Gromov–Hausdorff approximation $X\to Y$, some of its topological features are dominated by the corresponding ones of $Y$ (see for example [60]). Lemmas 4 and 4 below are in the spirit of this phenomenon.
[…]
For each $m\in\mathbb{N}$, there is $\varepsilon>0$ such that the following holds. Let $X$, $Y$ be proper geodesic spaces and maps $f:X\to Y$, $h:Y\to X$ such that
  1. (1)
    For each $r\in(0,1]$, $x\in X$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
  2. (2)
    For all $y_{1},y_{2}\in Y$ with $d(y_{1},y_{2})\leq 1$, one has $d(hy_{1},hy_{2})\leq d(y_{1},y_{2})+\varepsilon$.
  3. (3)
    For all $x\in X$, one has $d(h(f(x)),x)\leq\varepsilon$.
  4. (4)
    $f$ is continuous.
Then the map $f_{\ast}:H_{m}(X)\to H_{m}(Y)$ is injective. In particular, if $H_{m}(Y)$ is trivial, then so is $H_{m}(X)$.
Ideally, we would like the proof of Lemma 4 to go like this: $h\circ f\approx\text{Id}_{X}$, thus the identity map $H_{m}(X)\to H_{m}(X)$ factors as
$H_{m}(X)\xrightarrow{f_{\ast}}H_{m}(Y)\xrightarrow{h_{\ast}}H_{m}(X).$
This argument fails as $h$ is not necessarily continuous, so $h\circ f$ is not exactly homotopic to the identity on $X$. However, condition (1) allows us to get around these issues. We now turn to the actual proof (cf. [75, Proposition 1.3]).
Proof.
Assume $\varepsilon<\frac{1}{4m+3}$. For each $x_{0},\ldots,x_{k}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{k}\})\leq 3\varepsilon$ and $k\leq m+1$, we will construct a singular $k$-simplex $x_{0},\ldots,x_{k}]$ in $X$. For each $x_{0}\in X$, we denote by $x_{0}]$ the singular 0-simplex consisting of the point $x_{0}$. For each $x_{0},x_{1}\in X$ with $d(x_{0},x_{1})\leq 3\varepsilon,$ let $x_{0},x_{1}]$ be a geodesic from $x_{0}$ to $x_{1}$ seen as a singular 1-simplex.
Fix $k\leq m$ and assume by induction that for each $j\in\{1,\ldots,k\}$, and any points $x_{0},\ldots,x_{j}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{j}\})\leq 3\varepsilon$, there is a singular $j$-simplex $x_{0},\ldots,x_{j}]$ in $X$ contained in $B_{4j\varepsilon}(x_{0})$ and compatible with the lower dimensional simplices; that is,
(4.6) $\partial[x_{0},\ldots,x_{j}]=\sum_{\ell=0}^{j}\,(-1)^{\ell}\,[x_{0},\ldots,\hat{x}_{\ell},\ldots,x_{j}].$
Take $x_{0},\ldots,x_{k+1}\in X$ with ${\text{diam}}(\{x_{0},\ldots,x_{k+1}\})\leq 3\varepsilon$. Then the already constructed simplices $x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}]$ are contained in $B_{(4k+3)\varepsilon}(x_{0})$ and by (4.6) they satisfy
$\sum_{j=0}^{k+1}\,(-1)^{j}\,\partial\,[x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}]=0$
so by condition (1) there is a singular $(k+1)$-simplex $x_{0},\ldots,x_{k+1}]$ contained in $B_{4(k+1)\varepsilon}(x_{0})$ that restricted to its facets gives the ones already constructed; that is,
(4.7) $\partial[x_{0},\ldots,x_{k+1}]=\sum_{j=0}^{k+1}\,(-1)^{j}\,[x_{0},\ldots,\hat{x}_{j},\ldots,x_{k+1}].$
For each singular $k$-simplex $\tau:\Delta^{k}\to Y$, we define $h_{\sharp}(\tau)$ to be the singular $k$-simplex in $X$ given by $h\tau(e_{0}),\ldots,h\tau(e_{k})]$, where $e_{0},\ldots,e_{k}$ are the vertices of $\Delta^{k}$. By condition (2) this can be done provided ${\text{diam}}(\tau(\Delta^{k}))\leq\varepsilon$ and $k\leq m+1$. This definition expands linearly to all chains of the form $\psi=\sum_{\tau\in T}n_{\tau}\tau$ where $T$ is a finite set of singular $k$-simplices of diameter $\leq\varepsilon$ and $n_{\tau}\in\mathbb{Z}$ for each $\tau$, as
$h_{\sharp}(\psi):=\sum_{\tau\in T}n_{\tau}h_{\sharp}(\tau).$
Notice that by (4.7), we also have $\partial h_{\sharp}(\psi)=h_{\sharp}(\partial\psi)$.
Consider an $m$-cycle $\varphi=\sum_{\sigma\in S}n_{\sigma}\sigma$ in $X$ representing a homology class in ${\text{Ker}}(f_{\ast})$, where $S$ is a finite set of singular $m$-simplices, and $n_{\sigma}\in\mathbb{Z}$ for each $\sigma\in S$. Our goal is to show that $\varphi$ is a boundary. By (4), the cycle
$f_{\sharp}(\varphi)=\sum_{\sigma\in S}n_{\sigma}f_{\sharp}(\sigma):=\sum_{\sigma\in S}n_{\sigma}(f\circ\sigma)$
is well defined and by hypothesis is a boundary. Take $\psi$ a singular $(m+1)$-chain in $Y$ with $\partial\psi=f_{\sharp}(\varphi)$. By applying barycentric subdivision simultaneously on both sides, we can assume the diameter of any singular simplex featured in the chains $\varphi$ and $\psi$ has diameter $\leq\varepsilon$, so $h_{\sharp}(\psi)$ is well defined.
Then by condition (3), the boundary
$\partial h_{\sharp}(\psi)=h_{\sharp}(\partial\psi)=h_{\sharp}f_{\sharp}(\varphi)=\sum_{\sigma\in S}n_{\sigma}h_{\sharp}f_{\sharp}(\sigma)$
is uniformly close to $\varphi$. That is, for each $\sigma\in S$, the maps $\sigma,\,h_{\sharp}f_{\sharp}(\sigma):\Delta^{m}\to X$ are at uniform distance $\leq 8m\varepsilon.$ Thus if $\varepsilon$ is small enough (depending on $m$) one can (inductively on the skeleta just like at the beginning of the proof) construct a homotopy between $\varphi$ and $h_{\sharp}f_{\sharp}(\varphi)$. This implies that $\varphi$ is also a boundary.
[…]
Let $X$, $Y$ be proper geodesic spaces and maps $f:X\to Y$, $h:Y\to X$ that satisfy conditions (1), (2), (3), (4) from Lemma 4 with $\varepsilon=1/100$. Then $f_{\ast}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective. In particular, if $Y$ is simply connected then so is $X$.
Proof.
The proof is essentially the same as for Lemma 4 (cf. [64, Theorem 2.1]). Given a loop $\varphi:\mathbb{S}^{1}\to X$ representing an element of ${\text{Ker}}(f_{\ast})$, there is a disk $\psi:\mathbb{D}\to Y$ with $\psi|_{\mathbb{S}^{1}}=f\circ\varphi$. Triangulate $\mathbb{D}$ in such a way that the image of each face of the triangulation under $\psi$ has diameter $\leq 1/100$. Then for each face with vertices $a,b,c\in\mathbb{D}$, construct the singular 2-simplex $h(\psi(a)),h(\psi(b)),h(\psi(c))]$ in $X$ as in the proof of Lemma 4. Putting these simplices together, we obtain a continuous map $h_{\sharp}\psi:\mathbb{D}\to X$ with $h_{\sharp}\psi|_{\mathbb{S}^{1}}$ uniformly close to $\varphi$. Applying condition (1) once more we obtain that $\varphi$ is homotopic to $h_{\sharp}\psi|_{\mathbb{S}^{1}}$, hence nullhomotopic. ∎
[…]
At first, condition (4) in Lemma 4 may seem like an unnecessary technicality. However, simple examples such as Berger metrics on $\mathbb{S}^{3}$ converging to a round metric on $\mathbb{S}^{2}$ show that this condition is actually necessary.
[…]
With essentially the same proof, condition (1) in lemmas 4 and 4 can be replaced by the existence of a contractibility function in the sense of [60] (on which $\varepsilon$ depends). Since we will not need such version, we leave it as an exercise for the reader.
The following lemma is a basic result of algebraic topology.
[…]
Let $X$ be a simply connected topological manifold of dimension $n$. If $H_{m}(X)=0$ for all $m\in\{1,\ldots,n\}$, then $X$ is contractible.
Proof.
By [33, Proposition 3.29] all homology groups of $X$ vanish, so by Hurewicz Theorem [41, Theorem 2.1] and induction all homotopy groups vanish as well. By the work of Milnor [50, Corollary 1], $X$ is homotopy equivalent to a CW-complex and Whitehead Theorem [33, Theorem 4.5] applies, so $X$ is contractible. ∎
As mentioned in Section 1.1, in order to prove that a closed manifold is homeomorphic to an infranilmanifold, it is enough to verify that it is aspherical and it has the correct fundamental group. The following argument was extracted from [43].
[…]
Let $X$ be a closed aspherical manifold with $\pi_{1}(X)$ virtually nilpotent. Then $X$ is homeomorphic to an infranilmanifold.
Proof.
By the work of Lee–Raymond [47], there is an infranilmanifold $M$ with $\pi_{1}(M)=\pi_{1}(X)$. Then from the proof of the geometrization conjecture by Perelman in dimension $3$, the work of Freedman–Quinn in dimension $4$ [24], and the work of Farrell–Hsiang in dimensions 5 and higher [20], the Borel conjecture holds for such groups, thus the manifolds $X$ and $M$ are homeomorphic. ∎
5. Asphericity Theorem
As mentioned in Section 1.1, a crucial step to prove Theorem A is to show that in case of equality, the universal cover is a contractible manifold. We now prove something slightly stronger.
[…]
For each $K\in\mathbb{R}$, $N\geq 1$, there is $\varepsilon>0$ such that if $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space and $G\leq\text{Iso}(X)$ is a discrete group of isometries with ${\text{diam}}(X/G)\leq\varepsilon$, then
(5.1) ${\text{rank}}(G)\leq N,$
and in case of equality $X$ is homeomorphic to $\mathbb{R}^{N}$.
To prove Theorem 5 we follow the same strategy Gromov used to prove the Almost Flat Manifolds Theorem [30] (see [10] for a much more detailed explanation). After re-scaling, we can assume that $K$ is close to 0. Relying on results from [7, 76], we show that if $\varepsilon$ is small enough, then $X$ is Gromov–Hausdorff close to a nilpotent Lie group of dimension $\geq{\text{rank}}(G)$, proving (5.1).
In case ${\text{rank}}(G)=N$, we show that a finite-index subgroup $\Gamma\leq G$ is isomorphic to a lattice in a simply connected nilpotent Lie group $Y$. Furthermore, a large ball $B$ in $X$ is both homeomorphic and Gromov–Hausdorff close to a ball in the Euclidean space $\mathbb{R}^{N}$, the same being true for a ball $B^{\prime}$ in $Y$ when equipped with an appropriate invariant Riemannian metric. Moreover, the homeomorphism $f:B\to B^{\prime}$ can be taken to be almost equivariant with respect to the local actions of the group $\Gamma$ on both balls.
Using Lemma 4, the map $f$ is extended to a global $\Gamma$-equivariant homotopy equivalence $\tilde{f}:X\to Y$. Since the Borel Conjecture holds for virtually nilpotent groups, this finishes the proof. We now give the complete proof.
Proof.
Assuming the theorem fails, there is a sequence $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ of $\text{RCD}(K,N)$ spaces and $G_{i}\leq\text{Iso}(X_{i})$ discrete groups of isometries with ${\text{diam}}(X_{i}/G_{i})\to 0$ such that for all $i$ either ${\text{rank}}(G_{i})>N$, or ${\text{rank}}(G_{i})=N$ and $X_{i}$ is not homeomorphic to $\mathbb{R}^{N}$. After taking a subsequence and re-scaling the spaces $X_{i}$ by factors $\lambda_{i}$ with $\lambda_{i}\to\infty$ and $\lambda_{i}\cdot{\text{diam}}(X_{i}/G_{i})\to 0$, we can assume $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ is an $\text{RCD}(-\frac{1}{i},N)$ space for each $i$.
For $p_{i}\in X_{i}$, by taking again a subsequence and renormalizing the measures $\mathfrak{m}_{i}$, we can assume $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i},p_{i})$ converges in the pointed measured Gromov–Hausdorff sense to a pointed $\text{RCD}(K,N)$ space $(X,\mathsf{d},\mathfrak{m},p)$. By Lemma 3.10, there is $s\in\mathbb{N}$ and groups $G_{i}^{\prime}\leq G_{i}$ satisfying (3.2) and (3.3). By [76, Theorem 4.1], $X$ is a nilpotent Lie group equipped with an invariant metric. By Lemma 3.10 and Remark 3.10, for $i$ large enough we have
(5.2) ${\text{rank}}(G_{i})={\text{rank}}(G_{i}^{\prime})\leq n\leq N,$
where $n$ denotes the topological dimension of $X$. This finishes the proof of (5.1).
From now on we assume ${\text{rank}}(G_{i})=N$ for all $i$. Under this assumption, (5.2) implies that $X$ has topological dimension $N$. By Theorem 3.5, $\mathfrak{m}$ is a multiple of $\mathcal{H}^{N}$ so it is $\text{Iso}(X)$-invariant, hence by the work of Honda–Nepechiy [34, Theorem 2] the metric of $X$ is Riemannian. Moreover, by [51, Theorem 2.4], $X$ is abelian, and by Remark 3.10 we have $(X,p)=(\mathbb{R}^{N},0)$.
By Corollary 3.5, the groups $G_{i}^{\prime}$ admit no small subgroups. Then by Lemma 3.10 and Remark 3.10, for $i$ large enough there are elements $u_{1,i},\ldots,u_{n,i}\in\Gamma_{i}=G_{i}^{\prime}$, and $N_{1,i},\ldots,N_{n,i}\in\mathbb{R}^{+}$ with the following properties:
  • The set map $\alpha_{i}:\mathbb{Z}^{n}\to\Gamma_{i}$ given by
    $\alpha_{i}(m_{1},\ldots,m_{n})=u_{1,i}^{m_{1}}\cdots u_{n,i}^{m_{n}}$
    is a bijection. Moreover, there are polynomials $Q_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ of degree $\leq d(n)$ such that the product in $\Gamma_{i}$ is given by
    $g_{1}\cdot g_{2}=\alpha_{i}(Q_{i}(\alpha_{i}^{-1}(g_{1}),\alpha_{i}^{-1}(g_{2}))).$
  • There is $C>0$ such that the set
    $P_{i}:=P(u_{1,i},\ldots,u_{n,i};N_{1,i},\ldots,N_{n,i})\subset\Gamma_{i}$
    is a nilprogression in $C$-normal form with $\text{thick}(P_{i})\to\infty$.
  • For each $\varepsilon>0$ there is $\delta>0$ such that
    (5.3) $G(\delta P_{i})\subset\{$ $g\in\Gamma_{i}\,|\,\mathsf{d}_{i}(gp_{i},p_{i})\leq\varepsilon\},$
    (5.4) $\{g\in\Gamma_{i}\,|\,\mathsf{d}_{i}($ $gp_{i},p_{i})\leq\delta\}\subset G(\varepsilon P_{i})$
    for $i$ large enough.
By (5.4), after re-scaling the sequence $X_{i}$ by a fixed factor, we can assume
(5.5) $\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq 100\}\subset G(P_{i}).$
By (5.3), there is $\delta_{1}>0$ with $G(\delta_{1}P_{i})\subset\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq 1\}$. Hence there is $C_{1}\in\mathbb{N}$ so that
(5.6) $G(P_{i})\subset G(\delta_{1}P_{i})^{C_{1}}\subset\{g\in\Gamma_{i}|\mathsf{d}_{i}(gp_{i},p_{i})\leq C_{1}\}.$
Let $\mathfrak{l}_{i}$ be the Lie algebra of the Lie group $(\mathbb{R}^{n},Q_{i})$, and let
$\text{v}_{j,i}:=\log\left(u_{j,i}^{\lfloor\frac{N_{j,i}}{C}\rfloor}\right)\in\mathfrak{l}_{i}.$
By (5.6), after passing to a subsequence, for each $j\in\{1,\ldots,n\}$ we can assume
$u_{j,i}^{\lfloor\frac{N_{j,i}}{C}\rfloor}p_{i}\to\text{v}_{j}\,\,\text{ for some }\text{v}_{j}\in X=\mathbb{R}^{n}.$
By Lemmas 3.10 and 3.10, after further taking a subsequence, the groups $G_{i}^{\prime}$ converge equivariantly to the group of translations in $X=\mathbb{R}^{n}$. It is proven in [76, Lemma 8.2 and Lemma 2.64] that the structure coefficients of $\mathfrak{l}_{i}$ with respect to the basis $\{\text{v}_{1,i},\ldots,\text{v}_{n,i}\}$ converge to the structure coefficients in $\mathbb{R}^{n}$ with respect to $\{\text{v}_{1},\ldots,\text{v}_{n}\}$ as $i\to\infty$. These coefficients are of course 0 as $\mathbb{R}^{n}$ is abelian.
For each $i$, equip $\mathfrak{l}_{i}$ with the inner product $\langle\cdot,\cdot\rangle_{i}$ that makes the linear isomorphism $\mathfrak{l}_{i}\to\mathbb{R}^{n}$ given by $\text{v}_{j,i}\mapsto\text{v}_{j}$ an isometry and let $Y_{i}$ be the corresponding simply connected nilpotent Lie group equipped with the left invariant Riemannian metric induced from $\langle\cdot,\cdot\rangle_{i}$. Also let $q_{i}\in Y_{i}$ be the identity element and consider the map
$f_{i}:(X_{i},p_{i})\to(Y_{i},q_{i})$
obtained by composing the pointed Gromov–Hausdorff approximation $X_{i}\to X$ we have by definition, the isometric linear isomorphism $X\to\mathfrak{l}_{i}$ that sends $\text{v}_{j}\mapsto\text{v}_{j,i}$, and the exponential map $\mathfrak{l}_{i}\to Y_{i}$. Notice that by Theorem 3.6 we can assume $f_{i}|_{B_{100}(p_{i})}$ is continuous for $i$ large enough.
[…]
For each $\delta>0$, the following holds for $i$ large enough:
  1. (1)
    ${\text{diam}}(X_{i}/\Gamma_{i})$, ${\text{diam}}(Y_{i}/\Gamma_{i})\leq\delta$.
  2. (2)
    For all $x_{1},x_{2}\in B_{10}(p_{i})$,
    $|\textsf{d}(f_{i}(x_{1}),f_{i}(x_{2}))-\mathsf{d}_{i}(x_{1},x_{2})|\leq\delta.$
  3. (3)
    For all $y\in B_{1}(q_{i})$, there is $x\in B_{2}(p_{i})$ with
    $\textsf{d}(f_{i}(x),y)\leq\delta.$
  4. (4)
    If $x\in B_{10}(p_{i})$, $g\in\Gamma_{i}$ are such that $gf_{i}(x)\in B_{9}(q_{i})$, then
    $gx\in B_{10}(p_{i}).$
  5. (5)
    If $x\in B_{10}(p_{i})$, $g\in\Gamma_{i}$ are such that $gx\in B_{10}(p_{i})$, then
    $d(f_{i}(gx),g(f_{i}(x)))\leq\delta.$
Proof.
(1) is satisfied by hypothesis and the fact that the structure coefficients of $\mathfrak{l}_{i}$ converge to 0. (2) and (3) are satisfied as the maps $X_{i}\to X\to\mathfrak{l}_{i}\to Y_{i}$ are each either an isometry or a map uniformly close to being an isometry on a large ball.
By contradiction, if (5) doesn’t hold, after passing to a subsequence we could find $x_{i}\in B_{10}(p_{i}),$ $g_{i}\in\Gamma_{i}$ such that $g_{i}x_{i}\in B_{10}(p_{i})$ and
(5.7) $\textsf{d}(f_{i}(g_{i}x_{i}),g_{i}f_{i}(x_{i}))\geq\delta\,\text{ for all }i.$
By (5.5),
$g_{i}=\exp(t_{1,i}\text{v}_{1,i})\cdots\exp(t_{n,i}\text{v}_{n,i})$
for some $t_{1,i},\ldots,t_{n,i}\in[-1,1]$. After passing again to a subsequence we have $t_{j,i}\to t_{j}$ as $i\to\infty$ for some $t_{j}\in[-1,1]$. Then by how we defined the vectors $\text{v}_{j}$, it follows that $\exp(t_{j,i}\text{v}_{j,i})$ converges to the translation in $X$ by the vector $t_{j}\text{v}_{j}$. As the convergence is equivariant, $g_{i}$ converges to the translation by $t_{1}\text{v}_{1}+\ldots+t_{n}\text{v}_{n}$.
On the other hand, by looking at the pointed Gromov–Hausdorff approximations $Y_{i}\to X$ obtained by composing $\log:Y_{i}\to\mathfrak{l}_{i}$ with the linear isomorphism $\mathfrak{l}_{i}\to X$ that sends $\text{v}_{j,i}\mapsto\text{v}_{j}$, the sequence of actions of $\Gamma_{i}$ on $Y_{i}$ converge (again by Lemmas 3.10 and 3.10) to the group of translations of $X$. With respect to this convergence, by how we chose the map $\mathfrak{l}_{i}\to X$, the sequence $\exp(t_{j,i}\text{v}_{j,i})$ converges again to the translation in $X$ by $t_{j}\text{v}_{j}$, so $g_{i}$ again converges to $t_{1}\text{v}_{1}+\ldots+t_{n}\text{v}_{n}$.
The action of $g_{i}$ on both $X_{i}$ and $Y_{i}$ converges to the same isometry of $X$, and the pointed Gromov–Hausdorff approximations $f_{i}:X_{i}\to Y_{i}$ were obtained by composing the ones we used for $X_{i}\to X$ with the approximate inverses of the ones we used for $Y_{i}\to X$, meaning that $f_{i}$ approximately commutes with the action of $g_{i}$, contradicting (5.7).
Lastly, assume (4) fails, so after taking a subsequence there are $x_{i}\in B_{10}(p_{i})$, $g_{i}\in\Gamma_{i}$ with $g_{i}f_{i}(x_{i})\in B_{9}(q_{i})$ but $g_{i}x_{i}\not\in B_{10}(p_{i})$ for all $i$. After taking again a subsequence both sequences $x_{i}$ and $f_{i}(x_{i})$ converge to a point $x\in X$ and $g_{i}$ converges to an isometry of $X$. On one hand, since $g_{i}f_{i}(x_{i})\in B_{9}(q_{i})$ for all $i$, we have $gx\in B_{19/2}(p)$, but on the other hand since $g_{i}x_{i}\not\in B_{10}(p_{i})$ for all $i$ we also have $gx\not\in B_{19/2}(p)$. This is a contradiction. ∎
Now fix $\varepsilon>0$ given by Lemma 4 with $m=n$. By Proposition 5, the conditions of Lemma 4 hold for $i$ large enough so we have continuous $\Gamma_{i}$-equivariant maps $\tilde{f}_{i}:X_{i}\to Y_{i}$ such that
  1. (i)
    For all $x_{1},x_{2}\in X_{i}$ with $\min\{\mathsf{d}_{i}(x_{1},x_{2}),d(\tilde{f}_{i}(x_{1}),\tilde{f}_{i}(x_{2}))\}\leq 2$ one has
    $|\textsf{d}(\tilde{f}_{i}(x_{1})\tilde{f}_{i}(x_{2}))-\mathsf{d}_{i}(x_{1},x_{2})|\leq\varepsilon/3.$
  2. (ii)
    For all $y\in Y_{i}$, there is $x\in X_{i}$ with $\textsf{d}(\tilde{f}_{i}x,y)\leq\varepsilon/3$.
By (ii), we can define $h_{i}:Y_{i}\to X_{i}$ so that $\textsf{d}(\tilde{f}_{i}h_{i}(y),y)\leq\varepsilon/3$ for all $y\in Y_{i}$. Then from (i), follows that
  • For all $y_{1},y_{2}\in Y_{i}$ with $\textsf{d}(y_{1},y_{2})\leq 1$, one has $\mathsf{d}_{i}(h_{i}y_{1},h_{i}y_{2})\leq\textsf{d}(y_{1},y_{2})+\varepsilon$.
  • For all $x\in X_{i}$, one has $\mathsf{d}_{i}(h_{i}(\tilde{f}_{i}(x)),x)\leq\varepsilon$.
By Corollary 3.6, for $i$ large enough the following holds.
  • For each $r\in(0,1]$, $x\in X_{i}$, the ball $B_{r}(x)$ is contractible in $B_{r+\varepsilon}(x)$.
This means that all the conditions of Lemmas 4 and 4 are satisfied, hence the spaces $X_{i}$ are simply connected and have trivial homology up to degree $n$. By Corollary 3.6 the spaces $X_{i}$ are manifolds of dimension $n$ so from Lemma 4 they are contractible. Since the groups $\Gamma_{i}$ are discrete and torsion-free, they act freely and the quotient $X_{i}/\Gamma_{i}$ is a closed aspherical manifold. By Theorem 4, $X_{i}/\Gamma_{i}$ is homeomorphic to a nilmanifold, so its universal cover $X_{i}$ is homeomorphic to $\mathbb{R}^{n}=\mathbb{R}^{N}$. ∎
Proof of Theorem A.
By Theorem 5 applied to the universal covers, there is $\varepsilon(K,N)>0$ such that if ${\text{diam}}(X)\leq\varepsilon$ for an $\text{RCD}(K,N)$ space $X$, then ${\text{rank}}(\pi_{1}(X))\leq N$, and in case of equality $X$ is an aspherical manifold. From the proof of Theorem 5 we deduce that $\pi_{1}(X)$ is virtually nilpotent, so the result follows from Theorem 4. ∎
6. Eigenmap to a flat torus
Before we jump into the proof of Theorem B, we make a few remarks. If $(X,\mathsf{d},\mathfrak{m})$ is an $\text{RCD}(K,N)$ space, and $G\leq\text{Iso}(X)$ a group of measure preserving isometries, then $G$ admits a natural isometric linear action on $H^{1,2}(X)$ by
$g^{\ast}f(x):=f(g^{-1}x),\,\,\,\,g\in G,f\in H^{1,2}(X),x\in X.$
This action clearly commutes with the Laplacian. That is, if $f\in H^{1,2}(X)$ is in the domain of the Laplacian, then $g^{\ast}f$ is also in the domain of the Laplacian, and
$\Delta(g^{\ast}f)=g^{\ast}\Delta f.$
In particular, for each $\lambda\geq 0$, the eigenspace
$\{f\in H^{1,2}(X)|\Delta f=-\lambda f\}$
is a $G$-invariant subspace of $H^{1,2}(X)$.
Proof of Theorem B.
Assume by contradiction there is sequence of $\text{RCD}(K,N)$ spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ that converges in the measured Gromov–Hausdorff sense to $\mathbb{T}^{N}$ and there are abelian groups $G_{i}\leq\text{Iso}(X_{i})$ with ${\text{diam}}(X_{i}/G_{i})\to 0$, but the thesis of the theorem fails for each $X_{i}$.
By Lemma 3.10, the sequence $G_{i}$ converges equivariantly to a connected abelian group $G\leq\text{Iso}(\mathbb{T}^{N})$ acting freely and transitively, which can be naturally identified with $\mathbb{T}^{N}$. It is well known that the spectrum of the Laplacian on $\mathbb{T}^{N}$ consists of non-positive integers, the multiplicity of the eigenvalue 1-1 is $2N$, and the eigenfunctions corresponding to that eigenvalue are given by the coordinate functions. That is, if $\varphi_{j},\psi_{j}:\mathbb{T}^{N}\to\mathbb{R}$ are given by
(6.1) $x=(\varphi_{1}(x),\psi_{1}(x),\ldots,\varphi_{N}(x),\psi_{N}(x))\,\text{ for all }x\in\mathbb{T}^{N},$
then
$\Delta\varphi_{j}=-\varphi_{j},\,\Delta\psi_{j}=-\psi_{j}\,\text{ for each }j\in\{1,\ldots,N\},$
and $\{\varphi_{j},\psi_{j}\}_{j=1}^{N}$ is a set of orthogonal functions in $H^{1,2}(\mathbb{T}^{N})$ with the same norm. Set
$W_{j}:=\langle\varphi_{j}\rangle\oplus\langle\psi_{j}\rangle=\langle G\varphi_{j}\rangle,\hskip 85.35826ptV:=\bigoplus_{j=1}^{N}W_{j}.\hskip 34.14322pt$
Then it is not hard to check that for $f\in V$, one has
(6.2) $f\in W_{j}\text{ for some }j\hskip 14.22636pt\Longleftrightarrow\hskip 14.22636pt\dim_{\mathbb{R}}(\langle Gf\rangle)\leq 2.\hskip 36.98866pt$
For each $i\in\mathbb{N}$, set
$V_{i}:=\bigoplus_{1/2<\lambda<3/2}\{f\in H^{1,2}(X_{i})|\Delta f=-\lambda f.\}$
By Corollary 3.7, for $i$ large enough, $\dim_{\mathbb{R}}(V_{i})=2N$, and by Lemma 3.9,
(6.3) $V_{i}=\bigoplus_{j=1}^{N}W_{j,i}$
with each $W_{j,i}$ 2-dimensional, $G_{i}$-invariant, and spanned by eigenfunctions.
[…]
Assume for some choice of $j(i)\in\{1,\ldots,N\}$, a sequence of functions $f_{i}\in W_{j(i),i}$ converges in $H^{1,2}$ to some function $f\in H^{1,2}(\mathbb{T}^{N})$. Then $f\in W_{j_{0}}$ for some $j_{0}\in\{1,\ldots,N\}$.
Proof.
By (6.2), if the lemma fails, there are $n_{k}\in\mathbb{N}$ for $k\in\{1,2,3\}$, $\alpha_{k,\ell}\in\mathbb{R}$, $g_{k,\ell}\in G$ for $\ell\in\{1,\ldots,n_{k}\}$, such that the vectors
$\text{v}_{k}:=\sum_{\ell=1}^{n_{k}}\alpha_{k,\ell}g_{k,\ell}^{\,\ast}(f)\in V$
with $k\in\{1,2,3\}$ form an orthonormal set. Take elements $g_{k,\ell,i}\in G_{i}$ with $g_{k,\ell,i}\to g_{k,\ell}$ as $i\to\infty$. Then for fixed $k$, the vectors
$\text{v}_{k,i}:=\sum_{\ell=1}^{n_{k}}\alpha_{k,\ell}g_{k,\ell,i}^{\,\ast}(f_{i})\in W_{j(i),i}$
converge to $\text{v}_{k}$ in $H^{1,2}$. Hence as $i\to\infty$, for $k\neq\ell$ one has
$\|\text{v}_{k,i}\|_{H^{1,2}}\to 1,\hskip 85.35826pt\langle\text{v}_{k,i},\text{v}_{\ell,i}\rangle_{H^{1,2}}\to 0.$
This is impossible as $\dim_{\mathbb{R}}(W_{j(i),i})=2$. ∎
By Theorem 3.7(2), for each $j\in\{1,\ldots,N\}$ there are eigenfunctions $\varphi_{j,i}\in V_{i}$ with $\varphi_{j,i}\to\varphi_{j}$ uniformly and in $H^{1,2}$.
[…]
For each $j\in\{1,\ldots,N\}$, the quantity
(6.4) $\min\{\|\varphi_{j,i}-\text{v}\|_{H^{1,2}}|\,\text{v}\in W_{\ell,i},\ell\in\{1,\ldots,N\}\}$
goes to 0 as $i\to\infty$.
Proof.
If the lemma fails, after passing to a subsequence, there is a sequence $\ell(i)\in\{1,\ldots,N\}$ for which we can write
$\varphi_{j,i}=\varphi_{j,i}^{a}+\varphi_{j,i}^{b}$
with $\varphi_{j,i}^{a}\in W_{\ell(i),i}$, $\varphi_{j,i}^{b}\in W_{\ell(i),i}^{\perp}\leq V_{i}$, and
$\min\left\{\liminf_{i\to\infty}\|\varphi_{j,i}^{a}\|_{H^{1,2}},\liminf_{i\to\infty}\|\varphi_{j,i}^{b}\|_{H^{1,2}}\right\}>0.$
This is done by choosing $\ell(i)\in\{1,\ldots,N\}$ that attains the minimum in (6.4), and defining $\varphi_{j,i}^{a}$ to be the orthogonal projection of $\varphi_{j,i}$ to $W_{\ell(i),i}$. Also let $\text{w}_{i}\in W_{\ell(i),i}$ be a unit vector orthogonal to $\varphi_{j,i}^{a}$. Since each $\varphi_{j,i}^{a}$, $\varphi_{j,i}^{b}$, $\text{w}_{i}$ is a linear combination of eigenfunctions in $V_{i}$, by Theorem 3.7(1) after passing to a subsequence we can assume that
$\varphi_{j,i}^{a}\to\text{v}^{a},\hskip 28.45274pt\varphi_{j,i}^{b}\to\text{v}^{b},\hskip 28.45274pt\text{w}_{i}\to\text{w}$
uniformly and in $H^{1,2}$ for some mutually orthogonal non-zero vectors $\text{v}^{a},\text{v}^{b},\text{w}\in V$. By Lemma 6, $\{\text{v}^{a},\text{w}\}$ forms a basis of $W_{j_{0}}$ for some $j_{0}$, and $\text{v}^{b}\in W_{j_{0}}^{\perp}\leq V$. This would mean that $\varphi_{j}=\text{v}^{a}+\text{v}^{b}+\text{w}$ is not in $W_{j}$; a contradiction. ∎
Let $\varphi_{j,i}^{a}$, $\text{w}_{i}$ be as in the proof of Lemma 6 and let $\psi_{j,i}^{a}:=\|\varphi_{j,i}^{a}\|_{H^{1,2}}\text{w}_{i}\in W_{\ell(i),i}$. Then
$\varphi_{j,i}^{a}\to\varphi_{j},\hskip 85.35826pt\psi_{j,i}^{a}\to\pm\psi_{j}$
uniformly and in $H^{1,2}$ as $i\to\infty$. After possibly multiplying $\psi_{j,i}^{a}$ by 1-1, we can assume $\psi_{j,i}^{a}\to\psi_{j}$. We point out that the awkwardness of $\varphi_{j,i}^{a}$ being in $W_{\ell,i}$ instead of $W_{j,i}$ emanates from the fact that we have a repeated eigenvalue in the limit space, and we have a priori no way of telling apart the spaces $W_{j,i}$ (see the discussion before [13, Theorem 7.3]).
Consider the map $\Phi_{i}:X_{i}\to V_{i}$ given by
$\Phi_{i}(x)=\sum_{j=1}^{N}[\varphi_{j,i}^{a}(x)\varphi_{j,i}^{a}+\psi_{j,i}^{a}(x)\psi_{j,i}^{a}].$
Fix $g\in G_{i}$. Since $W_{j,i}$ is $G_{i}$-invariant for each $j$, one has
$g^{\ast}\varphi_{j,i}^{a}$ $=\cos(\theta_{j,i})\varphi_{j,i}^{a}+\sin(\theta_{j,i})\psi_{j,i}^{a},$
$g^{\ast}\psi_{j,i}^{a}$ $=-\sin(\theta_{j,i})\varphi_{j,i}^{a}+\cos(\theta_{j,i})\psi_{j,i}^{a}$
for some angles $\theta_{j,i}\in\mathbb{S}^{1}$. Hence
$g^{\ast}\Phi_{i}(x)$ $=\sum_{j=1}^{N}[g^{\ast}\varphi_{j,i}^{a}(gx)g^{\ast}\varphi_{j,i}^{a}+g^{\ast}\psi_{j,i}^{a}(gx)g^{\ast}\psi_{j,i}^{a}]$
$=\sum_{j=1}^{N}[\cos^{2}(\theta_{j,i})+\sin^{2}(\theta_{j,i})][\varphi_{j,i}^{a}(gx)\varphi_{j,i}^{a}+\psi_{j,i}^{a}(gx)\psi_{j,i}^{a}]$
$=\Phi_{i}(gx),$
so $\Phi_{i}$ is $G_{i}$-equivariant. For each $i$, let $U_{i}\subset V_{i}$ denote the set of vectors whose projection to $W_{j,i}$ is non-zero for all $j$, and let $\pi_{i}:U_{i}\to U_{i}$ be the map given by
$\pi_{i}(z_{1},\ldots,z_{N}):=\left(\frac{z_{1}}{\|z_{1}\|_{H^{1,2}}},\ldots,\frac{z_{N}}{\|z_{N}\|_{H^{1,2}}}\right)$
with $z_{j}\in W_{j,i}$ for each $j$. Since the decomposition in (6.3) is $G_{i}$-invariant, then so is the set $U_{i}$, and $\pi_{i}$ commutes with the action of $G_{i}$. Finally, we identify $V$ and $V_{i}$ with $\mathbb{R}^{2N}$ via the bases $\{\varphi_{1},\psi_{1},\ldots,\varphi_{N},\psi_{N}\}$ and $\{\varphi_{1,i}^{a},\psi_{1,i}^{a},\ldots,\varphi_{N,i}^{a},\psi_{N,i}^{a}\}$ respectively. After this identification, $\Phi_{i}$ converges uniformly to the isometric inclusion of $\mathbb{T}^{N}$ into $\mathbb{R}^{2N}$ given by (6.1). By Corollary 3.5 combined with Corollary 3.10, the action of $G_{i}$ on $X_{i}$ is free for large $i$. Then by Theorem 3.8 combined with Remark 3.8 and the fact that both $\Phi_{i}$ and $\pi_{i}$ are $G_{i}$-equivariant, the result follows. ∎
Proof of Theorem 1.
By contradiction, assume there is a sequence of $\text{RCD}(K,N)$ spaces $(X_{i},\mathsf{d}_{i},\mathfrak{m}_{i})$ with
${\text{diam}}(X_{i})\to 0,\hskip 85.35826pt\text{b}_{1}(X_{i})=N,$
but $X_{i}$ not bi-Hölder homeomorphic to a flat torus. In [52], it is shown that for $i$ large, $X_{i}$ is homeomorphic to $\mathbb{T}^{N}$, and a sequence of finite sheeted covers $X_{i}^{\prime}\to X_{i}$ is constructed in such a way that $X_{i}^{\prime}$ converges in the measured Gromov–Hausdorff sense to the flat torus $\mathbb{T}^{N}$.
Let $G_{i}\leq\text{Iso}(X_{i}^{\prime})$ be the group of deck transformations. By Theorem B, for $i$ large enough, there are free actions $G_{i}\to\text{Iso}(\mathbb{T}^{N})$ and $G_{i}$-equivariant bi-Hölder homeomorphisms $F_{i}:X_{i}^{\prime}\to\mathbb{T}^{N}$ that descend to bi-Hölder homeomorphisms $X_{i}=X_{i}^{\prime}/G_{i}\to\mathbb{T}^{N}/G_{i}$. Since $\mathbb{T}^{N}/G_{i}$ is locally isometric to $\mathbb{T}^{N}$, it is also a flat torus, which is a contradiction. ∎
Acknowledgments
The authors would like to thank Shouhei Honda, Raquel Perales, Xiaochun Rong, and Daniele Semola for helpful discussions, Vitali Kapovitch for sharing the reference [22], and Christine Escher for carefully reading and providing comments on a previous version of this paper. The authors are also grateful to two anonymous referees for their valuable feedback. While preparing this paper, Sergio Zamora held a Postdoctoral Fellowship at Max Planck Institute for Mathematics at Bonn. Xingyu Zhu held a Postdoctoral researcher position at the University of Bonn and gratefully acknowledge the financial support by the Deutsche Forschungsgemeinschaft (DFG) within the CRC 1060, at University of Bonn project number 211504053.